### v1

# The functional Loomis-Whitney type inequality in the Heisenberg groups and Projection theorems over finite fields

Abstract
We develop a functional Loomis-Whitney framework on the finite Heisenberg groups $\mathbb{H}^{n}(\mathbb{F}_{q})$ and discover connections to the boundedness and orthogonal projection problems. For $n=1$ we determine the sharp region of exponents $(u_{1},u_{2})$ for which the associated bilinear projection form is bounded uniformly in $q$, namely
$\frac{1}{u_{1}}+\frac{2}{u_{2}}\leq 2\quad\text{and}\quad\frac{2}{u_{1}}+\frac{1}{u_{2}}\leq 2,$
which includes the endpoint $L^{3/2}\times L^{3/2}\to L^{1}$. For general $n$ we prove a multilinear estimate at the critical exponent
$u=\frac{n(2n+1)}{n+1},$
via an induction on $n$ that exploits the group’s fiber structure together with multilinear interpolation. Specializing to indicators yields a sharp Loomis-Whitney type set inequality that controls $|K|$ for every finite $K\subset\mathbb{H}^{n}(\mathbb{F}_{q})$ by the sizes of its $2n$ Heisenberg projections $\{\pi_{j}(K)\}$, forcing a large projection in every configuration. A straightening map then converts these bounds into covering statements by additive cosets and provides a new approach to the orthogonal projection problem onto the vertical hyperplanes $\{x_{j}=0\}$, which presents an interesting link between commutative and non-commutative settings. The obtained results are optimal up to absolute constants, and in the planar case $n=1$, when the size of the set is not too small, our bound can be sharpened further using a point-line incidence estimate.
1 Introduction
The Loomis–Whitney inequality ([18]) bounds the volume of a $d$-dimensional body in $\mathbb{R}^{d}$ by the volumes of its (d1)(d-1)-dimensional coordinate projections. In the discrete setting it reads as follows. For $d\geq 2$, let $B_{1},\dots,B_{d}$ be finite sets, and let $A\subseteq B_{1}\times\cdots\times B_{d}$. For $i\in[d]$, define
$\pi_{i}\colon A\to B_{1}\times\cdots\times B_{i-1}\times B_{i+1}\times\cdots\times B_{d},\qquad\pi_{i}(a_{1},\ldots,a_{d}):=(a_{1},\ldots,a_{i-1},a_{i+1},\ldots,a_{d}).$
Then
$|A|^{\,d-1}\leq\prod_{i=1}^{d}|\pi_{i}(A)|.$ (1)
There are powerful entropy-theoretic generalizations of the inequality (1), notably Han’s inequality and Shearer’s inequality [19, 11]. Variants of this inequality, in relation to the set addition, also appear in additive combinatorics for abelian groups with numerous applications [21, 10].
This paper investigates such inequalities in the noncommutative setting, focusing on the Heisenberg group over a finite field. In this context, such inequalities are not only of independent interest but also shed light on group structures and related combinatorial problems. Some earlier works on the Heisenberg group can be found in [13, 14, 20, 15, 1]. Interestingly, this topic is connected to questions on boundedness and orthogonal projection problems. The approach developed here offers a new perspective on orthogonal projections onto coordinate planes. Further details will be discussed at the end of this section.
In the continuous setting, the Loomis–Whitney–type inequalities for the Heisenberg ( and more generally Carnot) groups have been studied extensively with applications in multilinear restriction theory and related areas; for example, see [8, 24, 3] and references therein.
We now recall the definition of the Heisenberg group. For a positive integer $n$, the $n$-th Heisenberg group $\mathbb{H}^{n}(\mathbb{F}_{q})$ is the set $\mathbb{F}^{2n+1}_{q}$ equipped with the group multiplication defined as follows: for $(x,t),(x^{\prime},t^{\prime})\in\mathbb{F}^{2n+1}_{q}$ with $x,x^{\prime}\in\mathbb{F}_{q}^{2n}$ and $t,t^{\prime}\in\mathbb{F}_{q},$
$(x,t)\cdot(x^{\prime},t^{\prime})=\left(x+x^{\prime},t+t^{\prime}+\frac{1}{2}\sum\limits_{j=1}^{n}(x_{j}x^{\prime}_{n+j}-x_{n+j}x^{\prime}_{j})\right).$ (2)
By definition, the group $\mathbb{H}^{n}(\mathbb{F}_{q})$ and the vector space $\mathbb{F}_{q}^{2n+1}$ have the same underlying sets. Thus, whenever necessary, we shall identify them only as sets: $\mathbb{H}^{n}(\mathbb{F}_{q})=\mathbb{F}_{q}^{2n+1}$. For each $1\leq j\leq 2n$, let
$W_{j}:=\{(x,t)\in\mathbb{F}_{q}^{2n+1}:x_{j}=0\},\qquad$ (3)
be the coordinate hyperplane corresponding to $x_{j}$, and let
$L_{j}:=\{(se_{j},0)\in\mathbb{F}_{q}^{2n+1}:s\in\mathbb{F}_{q}\}$
be the coordinate line corresponding to $x_{j}$, where $e_{j}$ ($j=1,\ldots,2n$) form the standard basis of $\mathbb{F}_{q}^{2n+1}.$
Then we easily see that $W_{j}$ and $L_{j}$ are subgroups of $\mathbb{H}^{n}(\mathbb{F}_{q})$ for $j=1,\ldots,2n$. Further, $L_{j}$ is the orthogonal complement $W_{j}^{\perp}$ of $W_{j}$, w.r.t the standard inner product on $\mathbb{F}_{q}^{2n+1}$; see Subsection 2.2 for these materials. The subgroups $W_{j}$ and $L_{j}$ will be often referred to as vertical and horizontal, respectively. We will provide more details on subgroup structures of $\mathbb{H}^{n}(\mathbb{F}_{q})$ in Section 2.
For $x\in\mathbb{F}^{2n}_{q}$ and $j\in\{1,\ldots,2n\}$ we use the notation $\widehat{x}_{j}$ to represent the element in $\mathbb{F}^{2n-1}_{q}$ obtained by entirely removing the $j$-th coordinate of $x$ while the notation $\check{x}_{j}$ to represent the element in $\mathbb{F}^{2n}_{q}$ obtained by setting the $j$-th coordinate of $x$ to 0. Let
$\begin{cases}\pi_{j}:\mathbb{H}^{n}(\mathbb{F}_{q})\rightarrow W_{j}\\ \pi_{n+j}:\mathbb{H}^{n}(\mathbb{F}_{q})\rightarrow W_{n+j}\end{cases}$ (4)
be the projections, respectively, defined by
$\begin{cases}\pi_{j}(x,t)=(\check{x}_{j},t+\frac{1}{2}x_{j}x_{n+j}),\\ \pi_{n+j}(x,t)=(\check{x}_{n+j},t-\frac{1}{2}x_{j}x_{n+j}).\end{cases}$ (5)
A direct computation shows that for each $j=1,\ldots,2n$, an element $(x,t)$ of $\mathbb{H}^{n}(\mathbb{F}_{q})$ can be written
$(x,t)=\pi_{j}(x,t)\cdot(x_{j}e_{j},0),$ (6)
where the dot denotes the group multiplication on $\mathbb{H}^{n}(\mathbb{F}_{q})$.
The fiber of $\pi_{j}$ at a point $(u,\tau)$ in $W_{j}$ is a line in the affine space $\mathbb{F}_{q}^{2n+1}$ passing through $(u,\tau)$. More precisely, for $j=1,\ldots,n$, the fiber of $\pi_{j}$ at $(u,\tau)\in W_{j}$ (so $u_{j}=0$) is given by
$\pi_{j}^{-1}\{(u,\tau)\}$ $=\big\{\,\big(u+se_{j},\;\tau-\tfrac{1}{2}\,u_{n+j}\,s\big):\;s\in\mathbb{F}_{q}\,\big\}=(u,\tau)\cdot L_{j},$
and the fiber of $\pi_{n+j}$ at $(u,\tau)\in W_{n+j}$ (so $u_{n+j}=0$) is given by
$\pi_{n+j}^{-1}\{(u,\tau)\}$ $=\big\{\,\big(u+se_{n+j},\;\tau+\tfrac{1}{2}\,u_{j}\,s\big):\;s\in\mathbb{F}_{q}\,\big\}=(u,\tau)\cdot L_{n+j}.$
That is, for each $j=1,\ldots,2n,$ the fibers of $\pi_{j}$ are exactly cosets of $W_{j}^{\perp}=L_{j}$.
Given a set $K$ in $\mathbb{H}^{n}(\mathbb{F}_{q})$, in this paper, we are interested in the Loomis–Whitney type inequality associated with $K$, namely, bounding the size of $K$ in terms of the sizes of its projections, and its generalizations.
For any set $S$ (in most cases, $S=\mathbb{F}_{q}^{2n})$, denote by $\mathscr{F}(S)$ the set of all $\mathbb{R}$-valued functions on $S$. Note that $\mathscr{F}(S)$ is a vector space over $\mathbb{R}.$ For each $1\leq r\leq\infty$, the $L^{r}$-norm on $\mathscr{F}(\mathbb{F}_{q}^{2n})$ is defined as
$\|f\|_{L^{r}(\mathbb{F}_{q}^{2n},dx)}=\left(\frac{1}{q^{2n}}\sum_{x\in\mathbb{F}_{q}^{2n}}|f(x)|^{r}\right)^{\frac{1}{r}}$
for $1\leq r<\infty$, and for $r=\infty$
$\|f\|_{L^{\infty}(\mathbb{F}_{q}^{2n},dx)}=\max\limits_{x\in\mathbb{F}_{q}^{2n}}|f(x)|.$
Throughout, for each $j=1,\ldots,2n,$ we shall identify $W_{j}=\mathbb{F}_{q}^{2n}$ via the map $\phi_{j}:W_{j}\stackrel{{\scriptstyle\sim}}{{\rightarrow}}\mathbb{F}_{q}^{2n}$ defined by $(\check{x}_{j},t)\mapsto(\widehat{x}_{j},t)$. For simplicity, for a function $f$ on $\mathbb{F}_{q}^{2n}$ and $(x,t)\in\mathbb{H}^{n}(\mathbb{F}_{q})$, we write $f(\pi_{j}(x,t))$ for $f(\phi_{j}(\pi_{j}(x,t)))$.
Definition 1.1.
Let $n$ be a positive integer. For extended real numbers $1\leq u_{1},u_{2},\ldots,u_{2n}\leq\infty$, define $LW(u_{1},u_{2},\ldots,u_{2n})$ as the smallest constant, if it exists, such that the inequality
$\frac{1}{q^{2n+1}}\sum_{(x,t)\in\mathbb{H}^{n}(\mathbb{F}_{q})}\prod_{j=1}^{2n}f_{j}(\pi_{j}(x,t))\leq LW(u_{1},u_{2},\ldots,u_{2n})\prod_{j=1}^{2n}\|f_{j}\|_{L^{u_{j}}(\mathbb{F}_{q}^{2n},dx)}$ (7)
holds for all nonnegative $f_{1},f_{2},\ldots,f_{2n}\in\mathscr{F}(\mathbb{F}_{q}^{2n})$. If it does not exist, then we set $LW(u_{1},u_{2},\ldots,u_{2n}):=\infty.$
When $f_{1}=\cdots=f_{2n}$ are all the indicator functions of a subset $K$ of $\mathbb{H}^{n}(\mathbb{F}_{q})$, then the above inequality presents the Loomis-Whitney type inequality.
We first present a result in the case $n=1$.
Theorem 1.2.
For $1\leq u_{1},u_{2}\leq\infty,$ the following holds.
$LW(u_{1},u_{2})\lesssim 1\quad\mbox{if and only if}\quad\frac{1}{u_{1}}+\frac{2}{u_{2}}\leq 2\quad\mathrm{and}\quad\frac{2}{u_{1}}+\frac{1}{u_{2}}\leq 2.$
In this article we write $X\lesssim Y$ to mean that there exists some constant $C>0$ such that $X\leq CY$.
Corollary 1.3.
For any nonnegative $f_{1},f_{2}\in\mathscr{F}(\mathbb{F}_{q}^{2})$, we have
$\frac{1}{q^{3}}\sum_{(x,t)\in\mathbb{H}^{1}(\mathbb{F}_{q})}f_{1}(\pi_{1}(x,t))f_{2}(\pi_{2}(x,t))\lesssim\|f_{1}\|_{L^{3/2}(\mathbb{F}_{q}^{2},dx)}\|f_{2}\|_{L^{3/2}(\mathbb{F}_{q}^{2},dx)}.$
When $n>1$, it is difficult to find necessary and sufficient conditions on $u_{1},u_{2},\ldots,u_{2n}$ under which $LW(u_{1},u_{2},\ldots,u_{2n})\lesssim 1$, due to the lack of understanding of structures of the group $\mathbb{H}^{n}(\mathbb{F}_{q})$. However, by using Corollary 1.3 and the inductive argument introduced in the continuous case [8], we are able to prove the following generalization of Corollary 1.3 for $u_{1}=\cdots=u_{2n}=\frac{n(2n+1)}{n+1}$.
Theorem 1.4.
For any nonnegative $f_{1},\ldots,f_{2n}\in\mathscr{F}(\mathbb{F}_{q}^{2n})$, we have
$\frac{1}{q^{2n+1}}\sum_{(x,t)\in\mathbb{H}^{n}(\mathbb{F}_{q})}\prod_{j=1}^{2n}f_{j}(\pi_{j}(x,t))\lesssim\prod_{j=1}^{2n}\|f_{j}\|_{L^{\frac{n(2n+1)}{n+1}}(\mathbb{F}_{q}^{2n},dx)}.$ (8)
As an application of Theorem 1.4, we obtain the following.
Corollary 1.5.
For $K\subset\mathbb{H}^{n}(\mathbb{F}_{q})$, we have
$|K|\lesssim q^{\frac{1}{2n+1}}\prod_{j=1}^{2n}|\pi_{j}(K)|^{\frac{n+1}{n(2n+1)}}.$ (9)
In particular, for $K\subset\mathbb{H}^{1}(\mathbb{F}_{q})$,
$\max\{|\pi_{1}(K)|,|\pi_{2}(K)|\}\gtrsim|K|^{\frac{3}{4}}q^{-\frac{1}{4}}.$
Without making additional assumptions, we can easily check that the upper bound for $|K|$ in (9) is sharp. Indeed, for $n=1,$ fix $t_{0}\in\mathbb{F}_{q}$ and let $K:=\{(x_{1},0,t_{0})\colon x_{1}\in\mathbb{F}_{q}\}$. Then we have $|K|=q$. Moreover, $|\pi_{1}(K)|=1$ and $|\pi_{2}(K)|=q$. Thus, the bound $|K|\leq q^{\frac{1}{3}}|\pi_{1}(K)|^{\frac{2}{3}}|\pi_{2}(K)|^{\frac{2}{3}}$ is sharp.
In light of this example, we cannot replace the factor $q^{\frac{1}{2n+1}}$ by $|\pi_{v}(K)|^{\frac{1}{2n+1}}$, where $\pi_{v}(K)$ is the image of $K$ under the (vertical) projection $\pi_{v}:\mathbb{H}^{n}(\mathbb{F}_{q})\rightarrow\mathbb{F}_{q}$ defined by $(x,t)\mapsto t$. This example can be extended to larger $n$ by setting
$K\;=\;\Big\{(x,t_{0})\in\mathbb{H}^{n}(\mathbb{F}_{q}):\ x_{n+1}=\cdots=x_{2n}=0\Big\}.$
Then $|K|=q^{n}$, and for the vertical projections $\pi_{1},\dots,\pi_{2n}$ we have
$|\pi_{j}(K)|=\begin{cases}q^{n-1},&1\leq j\leq n,\\[2.0pt] q^{\,n},&n+1\leq j\leq 2n.\end{cases}$
Consequently,
$q^{n}\;=\;|K|=q^{\frac{1}{2n+1}}\prod_{j=1}^{2n}|\pi_{j}(K)|^{\frac{n+1}{n(2n+1)}}.$
When $n=1$ and $|K|>q$, by using the point-line incidence theorem due to Vinh [23], we are able to provide a better lower bound for the quantity $\max\{|\pi_{1}(K)|,|\pi_{2}(K)|\}$.
Theorem 1.6.
For $K\subset\mathbb{H}^{1}(\mathbb{F}_{q})$, we have
$|K|\lesssim\frac{|\pi_{1}(K)|\,|\pi_{2}(K)|}{q}+q^{\frac{1}{2}}\sqrt{|\pi_{1}(K)|\,|\pi_{2}(K)|}.$
In particular,
$\max\{|\pi_{1}(K)|,|\pi_{2}(K)|\}\gtrsim\min\left\{|K|^{\frac{1}{2}}q^{\frac{1}{2}},~|K|q^{-\frac{1}{2}}\right\}.$
This theorem is also sharp in its range. To see this, let $A,B\subset\mathbb{F}_{q}$ with $|A|=|B|=m$, and put $T=\mathbb{F}_{q}$. Define
$K:=A\times B\times T\;\subset\;\mathbb{H}^{1}(\mathbb{F}_{q}).$
Then
$\pi_{1}(K)=B\times\mathbb{F}_{q},\quad|\pi_{1}(K)|=mq,\qquad\pi_{2}(K)=A\times\mathbb{F}_{q},\quad|\pi_{2}(K)|=mq,$
and
$|K|=|A|\,|B|\,|T|=m^{2}q.$
Consequently,
$\frac{|\pi_{1}(K)|\,|\pi_{2}(K)|}{q}\;+\;q^{1/2}\sqrt{|\pi_{1}(K)|\,|\pi_{2}(K)|}\;=\;m^{2}q\;+\;mq^{3/2},$
which matches $|K|$ up to absolute constants.
Now we address how our results are connected to some questions in the literature.
Bhowmik, Iosevich, Koh, and Pham’s boundedness problem:
Let $\mathcal{K}\colon\mathbb{F}_{q}^{d}\times\mathbb{F}_{q}^{d}\to\mathbb{R}$ be any function (which, below, will play a role of the kernel of an integral operator for the counting measure), and $\mathcal{G}=(V,E)$ a connected ordered graph on $n$ vertices. In a recent paper [4], Bhowmik, Iosevich, Koh, and Pham introduced the boundedness problem for a multilinear operator associated with $\mathcal{K}$ and $\mathcal{G}$. To be precise, define
$\Lambda=\Lambda_{\mathcal{G}}^{\mathcal{K}}:\mathscr{F}(\mathbb{F}_{q}^{d})\times\cdots\times\mathscr{F}(\mathbb{F}_{q}^{d})\rightarrow\mathbb{R}$
by
$(f_{1},\ldots,f_{m})\mapsto\frac{1}{\mathcal{N}(\mathcal{G})}\sum_{x^{1},\dots,x^{m}\in{\mathbb{F}}_{q}^{d}}\ \prod_{(i,j)\in{E}(\mathcal{G})}\mathcal{K}(x^{i},x^{j})\prod_{l=1}^{m}f_{l}(x^{l}),$
where $\mathcal{N}(\mathcal{G})$ is the number of distinct embeddings of $\mathcal{G}$ in $\mathbb{F}_{q}^{d}$. The boundedness problem for operator $\Lambda_{\mathcal{G}}^{\mathcal{K}}$ asks if one can determine numbers $1\leq u_{1},\ldots,u_{m}\leq\infty$ for which there is a (smallest) constant $\Lambda_{\mathcal{G}}^{\mathcal{K}}(u_{1},\dots,u_{m})>0$ such that the inequality
$\frac{1}{\mathcal{N}(\mathcal{G})}\sum_{x^{1},\dots,x^{m}\in{\mathbb{F}}_{q}^{d}}\ \prod_{(i,j)\in{E}(\mathcal{G})}\mathcal{K}(x^{i},x^{j})\prod_{l=1}^{m}f_{l}(x^{l})\leq\Lambda_{\mathcal{G}}^{\mathcal{K}}(u_{1},\dots,u_{m})\prod_{i=1}^{m}{\|f_{i}\|}_{L^{u_{i}}(\mathbb{F}_{q}^{d},dx)}$ (10)
holds for all nonnegative $f_{1},\ldots,f_{m}\in\mathscr{F}(\mathbb{F}_{q}^{d})$. Using the discrete Fourier analysis, they settled the case where $\mathcal{G}$ is a connected graph with at most $4$ vertices and $\mathcal{K}(x,y)$ is defined by $\mathcal{K}(x,y)=1$ if $\|x-y\|=(x_{1}-y_{1})^{2}+(x_{2}-y_{2})^{2}=1$, and $\mathcal{K}(x,y)=0$ otherwise.
The operator $\Lambda_{\mathcal{G}}^{\mathcal{K}}$ is a general form of several geometric counting functions:
  1. 1.
    If $\mathcal{G}=K_{2}$, and $f_{1}=f_{2}$ are the indicator functions of a set $A$ in $\mathbb{F}_{q}^{d}$, then the operator $\Lambda_{\mathcal{G}}^{\mathcal{K}}$ counts the number of pairs of points in $A\times A$ of distance $1$.
  2. 2.
    If $\mathcal{G}=C_{4}$, and $f_{1}=f_{2}=f_{3}=f_{4}$ are the indicator functions of a set $A$ in $\mathbb{F}_{q}^{d}$, then the operator $\Lambda_{\mathcal{G}}^{\mathcal{K}}$ counts the number of rhombi of side-length $1$ in $A$.
  3. 3.
    If $\mathcal{G}$ is a tree and $f_{1}=\cdots=f_{m}$ are the indicator functions of a set $A$ in $\mathbb{F}_{q}^{d}$, then the operator $\Lambda_{\mathcal{G}}^{\mathcal{K}}$ counts the number of copies of the given tree $\mathcal{G}$ in $A$.
The inequality (7) for $n=1$ is essentially the same as the inequality (10) for $m=2$, with suitable $\mathcal{G}$ and $\mathcal{K}$ chosen. To see this, define $\phi\colon\mathbb{F}_{q}^{3}\to\mathbb{F}_{q}^{3}$ by $\phi(x,y,t)=(x,y,t+\frac{xy}{2})$. Then, since $\phi$ is a bijective map, the left hand side of (7) can be written
$\frac{1}{q^{3}}\sum_{(x_{1},x_{2},t)\in\mathbb{H}_{q}^{1}}f_{1}(\pi_{1}(x_{1},x_{2},t))f_{2}(\pi_{2}(x_{1},x_{2},t))=\frac{1}{q^{3}}\sum_{x_{1},x_{2},t\in\mathbb{F}_{q}}f_{1}(\pi_{1}(\phi(x_{1},x_{2},t)))f_{2}(\pi_{2}(\phi(x_{1},x_{2},t)))$
$=\frac{1}{q^{3}}\sum_{x_{1},x_{2},t\in\mathbb{F}_{q}}f_{1}(x_{2},t+x_{1}x_{2})f_{2}(x_{1},t).$ (11)
Now let $\mathcal{G}$ be the complete graph of two vertices and define $\mathcal{K}$ by
$\mathcal{K}((a_{1},a_{2}),(b_{1},b_{2}))=\begin{cases}1~&\mbox{if}~a_{1}b_{1}+b_{2}-a_{2}=0,\\ 0~&\mbox{otherwise}.\end{cases}$
Then the left hand side $\Lambda_{\mathcal{G}}^{\mathcal{K}}(f_{1},f_{2})$ of the inequality (10) is equal to
$\frac{1}{q^{3}}\sum_{(a_{1},a_{2}),(b_{1},b_{2})\in\mathbb{F}^{2}_{q},a_{1}b_{1}+b_{2}=a_{2}}f_{1}(a_{1},a_{2})f_{2}(b_{1},b_{2}),$
which coincides with the value (1). Thus, in this case, both inequalities (7) and (10) coincide with each other. This connection suggests a further study of the operator $\Lambda_{\mathcal{G}}^{\mathcal{K}}$ for various $\mathcal{K}$ and $\mathcal{G}$.
The orthogonal projection problem:
Fix a subset $K\subset\mathbb{F}_{q}^{2n+1}$. For $r\geq 1,$ let
[…]
where, by translates of $W^{\perp}$, we mean $\{x+W^{\perp}\colon x\in\mathbb{F}_{q}^{2n+1}\}$. Chen [6] proved that
$|\mathscr{E}_{r}|\;\leq\;\begin{cases}q^{2n-1}\,r,&\text{if }r\leq\dfrac{|K|}{2},\\[5.69054pt] \dfrac{r\,q^{4n}}{(q^{2n}-r)\,|K|},&\text{if }0<r<q^{2n}.\end{cases}$ (12)
(Informally, it states that for most subspaces $W\subset\mathbb{F}_{q}^{2n+1}$, the set $K$ cannot be covered by a relatively “small” number of translates of $W^{\perp}.$)
Let $\mathscr{H}=\{W_{1},\dots,W_{2n}\}$ be the set of the vertical hyperplanes of $\mathbb{H}^{n}(\mathbb{F}_{q})$, given in (3). If $|\mathscr{H}|>|\mathscr{E}_{r}|$, then $\mathscr{H}\nsubseteq\mathscr{E}_{r}$ and hence there exists $W_{j}\in\mathscr{H}$ such that $K$ cannot be covered by $\leq r$ translates of $W_{j}^{\perp}$.
Define $T_{j}:\mathbb{F}_{q}^{2n+1}\to\mathbb{F}_{q}^{2n+1}$ by
$T_{j}(x,t)=\begin{cases}(x,\ t+\tfrac{1}{2}\,x_{j}x_{n+j}),&1\leq j\leq n,\\[2.0pt] (x,\ t-\tfrac{1}{2}\,x_{j-n}\,x_{j}),&n+1\leq j\leq 2n.\end{cases}$ (13)
This map is well-defined and bijective. We observe that for every $j\in\{1,\dots,2n\}$ and every $(u,\tau)\in W_{j}$,
$T_{j}\big(\pi_{j}^{-1}\{(u,\tau)\}\big)=(u,\tau)+W_{j}^{\perp},$
i.e. $T_{j}$ maps each Heisenberg fiber bijectively onto an additive translate of $W_{j}^{\perp}$.
Notice that $|T_{j}(K)|=|K|$. Thus, the minimal number of translates of $W_{j}^{\perp}$ that cover $T_{j}(K)$ equals $|\pi_{j}(K)|$. Therefore
$|\mathscr{H}|>|\mathscr{E}_{r}|\quad\Longrightarrow\quad\exists\,j\in\{1,\dots,2n\}:\;|\pi_{j}(K)|>r.$ (14)
To describe the connection between some of our results and Chen’s, assume that
$r\;<\;\frac{n\,|K|}{q^{2n}}.$ (15)
Then from the second line of (12), we easily obtain $|\mathscr{E}_{r}|<2n=|\mathscr{H}|$. Thus, by (14) there exists some $j$ with $|\pi_{j}(K)|>r$.
Since this holds for every $r$ satisfying (15), we obtain
$\quad\max_{1\leq j\leq 2n}|\pi_{j}(K)|\;\gtrsim_{n}\;\frac{|K|}{q^{2n}}\quad$
(up to a constant depending only on $n$). This lower bound extracted from Chen’s result for the tiny family $\mathscr{H}$, is substantially weaker than our Heisenberg Loomis–Whitney bound $\max_{j}|\pi_{j}(K)|\;\gtrsim\;|K|^{\frac{2n+1}{2(n+1)}}\,q^{-\frac{1}{2(n+1)}},$ given in Corollary 1.5. To be precise,
$\frac{|K|^{\frac{2n+1}{2(n+1)}}q^{-\frac{1}{2(n+1)}}}{|K|/q^{2n}}\;=\;|K|^{-\frac{1}{2n+2}}\;q^{\,2n-\frac{1}{2(n+1)}}\;\geq\;\,q^{\,2n-1}\,.$
Since $T_{j}$ are bijective maps, the method developed in this paper offers a new approach to address the orthogonal projection problem on the family of vertical hyperplanes.
In light of these observations, it is natural to investigate projection problems in the Heisenberg group, paralleling the continuous theory of Balogh, Fassler, Mattila, and Tyson [2]. We will develop this in a sequel paper.
2 Subgroups of the Heisenberg group and projections
This section is devoted to the study of subgroups $G$ of the Heisenberg group $\mathbb{H}^{n}(\mathbb{F}_{q})$ and the analysis of the projections onto relevant subgroups. The structural statements developed here will be used throughout the paper and provide tools for subsequent work.
2.1 Basics on subgroups of $\mathbb{H}^{n}(\mathbb{F}_{q})$
By a group structure on a vector space $V$ over $\mathbb{F}_{q}$ we mean the underlying additive group structure on $V$, forgetting the scalar multiplication. Trivially, a subspace of the vector space $V$ is a subgroup of $V$. If the field $\mathbb{F}_{q}$ is a prime field, i.e., $q=p$ for some prime $p$, then the converse is also true. Indeed, if $W$ is a subgroup of $V,$ then we can give a scalar multiplication on $W$ that makes $W$ into a subspace of $V.$ To see this, we may assume $\mathbb{F}_{p}=\{0,1,\ldots,p-1\}.$ For $t\in\mathbb{F}_{p}$ and $w\in W$, define $tw:=w+\cdots+w$ ($t$-times). We easily see that this coincides with the scalar multiplication on $W$ induced from $V$, and so $W$ is a subspace of $V.$ If $\mathbb{F}_{q}$ is not a prime field, then there is a subgroup $W$ of $V$ that is not a subspace. For example, if $q=p^{r}$ for some $r\geq 2,$ the subfield $\mathbb{F}_{p}$ is a subgroup of the (one dimensional) vector space $\mathbb{F}_{q},$ but not a subspace of $\mathbb{F}_{q}$ (over $\mathbb{F}_{q}$). Over a general $\mathbb{F}_{q}$ with $q=p^{r},$ in order to obtain all subgroups of a vector space $V$, it is enough to view $V$ as a vector space over $\mathbb{F}_{p}$ and find all subspaces of $V$ over $\mathbb{F}_{p}.$ Similarly, linear maps between two vector spaces over a prime field are precisely group homomorphisms between the corresponding underlying groups.
The quadratic term in the group multiplication (2) on $\mathbb{H}^{n}(\mathbb{F}_{q})$ plays an important role in understanding $\mathbb{H}^{n}(\mathbb{F}_{q})$ and its subgroups. So we treat it especially. For $x,x^{\prime}\in\mathbb{F}_{q}^{2n}$, define
$\omega(x,x^{\prime}):=\sum_{i=1}^{n}(x_{i}x_{n+i}^{\prime}-x_{i}^{\prime}x_{n+i}).$
Then $\omega$ is a symplectic form, i.e., a nondegenerate skew-symmetric bilinear form on $\mathbb{F}_{q}^{2n}$. Note that $\omega$ is determined by the skew symmetric matrix $A_{n}$, i.e., $\omega(x,x^{\prime})=x\cdot A_{n}\cdot(x^{\prime})^{t},$ where
$A_{n}:=\begin{bmatrix}0&I_{n}\\ -I_{n}&0\\ \end{bmatrix}$,
and $(\,)^{t}$ stands for the transpose of a matrix.
Definition 2.1.
A subspace $S$ of $\mathbb{F}_{q}^{2n}$ is called isotropic if $\omega(x,x^{\prime})=0$ for any two vectors $x,x^{\prime}\in S$.
Note that by linear algebra, maximal isotropic subspaces of $\mathbb{F}_{q}^{2n}$ have dimension $n$, and all subspaces of dimension one are isotropic.
Let $\pi_{h}:\mathbb{H}^{n}(\mathbb{F}_{q})\rightarrow\mathbb{F}_{q}^{2n}$ be the projection defined by $\pi_{h}(x,t)=x$. Then $\pi_{h}$ is a group homomorphism and so if $G$ is a subgroup of $\mathbb{H}^{n}(\mathbb{F}_{q}),$ then the image $\pi_{h}(G)$ is a subgroup of $\mathbb{F}_{q}^{2n}.$ Recall that $\pi_{v}:\mathbb{H}^{n}(\mathbb{F}_{q})\rightarrow\mathbb{F}_{q}$ is the projection defined by $\pi_{v}(x,t)=t$. Note that $\pi_{v}$ is not a group homomorphism, but its restriction to the ‘vertical line’ $\{(0,t)\in\mathbb{H}^{n}(\mathbb{F}_{q})\,|\,t\in\mathbb{F}_{q}\}$ is a group isomorphism.
Define an action of the group $\mathbb{F}_{q}^{*}$ on $\mathbb{H}^{n}(\mathbb{F}_{q})$ by $s\cdot(x,t)=(sx,s^{2}t)$ for $s\in\mathbb{F}_{q}^{*}$ and $(x,y)\in\mathbb{H}^{n}.$ This action shall be called the dilation action, following [2].
Definition 2.2.
A subgroup $G$ of $\mathbb{H}^{n}$ is called homogeneous if $G$ is invariant under the dilation action of $\mathbb{F}_{q}^{*}$ on $\mathbb{H}^{n}$.
Recall that the orbit of an element $(x,t)\in\mathbb{H}^{n}$ under the dilation action is given by
$\mathcal{O}_{\mathbb{F}_{q}^{*}}(x,t):=\{s\cdot(x,t)\in\mathbb{H}^{n}(\mathbb{F}_{q})\,|\,s\in\mathbb{F}_{q}^{*}\}.$
If $G$ is any homogeneous subgroup of $\mathbb{H}^{n}$ containing $(x,t)$, then $G$ necessarily includes $\mathcal{O}_{\mathbb{F}_{q}^{*}}(x,t)$.
Lemma 2.3.
For $(x,t)\in\mathbb{H}^{n}(\mathbb{F}_{q})$, the cardinality of $\mathcal{O}_{\mathbb{F}_{q}^{*}}(x,t)$ is given as follows.
  • (i)
    If $x\neq 0$, then $|\mathcal{O}_{\mathbb{F}_{q}^{*}}(x,t)|=q-1.$
  • (ii)
    If $x=0$ and $t\neq 0$, then $|\mathcal{O}_{\mathbb{F}_{q}^{*}}(x,t)|=\frac{q-1}{2}.$
  • (iii)
    If $x=0$ and $t=0$, then $|\mathcal{O}_{\mathbb{F}_{q}^{*}}(x,t)|=1.$
Proof.
We can easily check that the isotropy groups of $(x,t)$ for the cases (i), (ii), (iii)
$\mathrm{ISO}_{\mathbb{F}_{q}^{*}}(x,t)=\{s\in\mathbb{F}_{q}^{*}\,|\,s\cdot(x,t)=(x,t)\}$
are $\{0\},$ $\mathbb{Z}_{2}$ and $\mathbb{F}_{q}^{*}$ , respectively. Then the lemma is immediate from the formula
$|\mathcal{O}_{\mathbb{F}_{q}^{*}}(x,t)||\mathrm{ISO}_{\mathbb{F}_{q}^{*}}(x,t)|=|\mathbb{F}_{q}^{*}|=q-1.$
Now we will identify and count all subgroups $G$ of $\mathbb{H}^{n}$ for two special cases, namely the case when $\mathbb{F}_{q}$ is a prime field, and the case where subgroups $G$ are homogeneous.
First assume that $\mathbb{F}_{q}$ is a prime field, say, $q=p$ for some prime $p.$ In this case we write $\mathbb{F}_{p}$ for $\mathbb{F}_{q}$. First fix a subgroup (equivalently, a subspace) $S$ of $\mathbb{F}_{p}^{2n}.$ If $G$ is a subgroup of $\mathbb{H}^{n}(\mathbb{F}_{p})$ such that $\pi_{h}(G)=S$, by cardinality reason for finite groups, there are only two possibilities
$|G|=|S|,$
$|G|=|S|p.$
If $|G|=|S|p,$ then necessarily $G=S\times\mathbb{F}_{p}.$ If $|G|=|S|,$ then $G$ is a graph $\Gamma_{S}(\rho)$ of a map $\rho:S\rightarrow\mathbb{F}_{p}$,
$\Gamma_{S}(\rho):=\{(x,t)\subset\mathbb{H}^{n}\,|\,x\in S,t=\rho(x)\}.$
But since $G=\Gamma_{S}(\rho)$ is a group, i.e., closed under the induced multiplication, $S$ should be isotropic and $\rho$ is a group homomorphism. Summarizing, we have
Proposition 2.4.
If $G$ is a subgroup of $\mathbb{H}^{n}(\mathbb{F}_{p})$ for a prime $p$, then $G$ is one of following two types
  • (a)
    $G=S\times\mathbb{F}_{p}$ for a subspace $S$ of $\mathbb{F}_{p}^{2n}$.
  • (b)
    $G=\Gamma_{S}(\rho)$ for a group homomorphism (equivalently, a linear map) $\rho:S\rightarrow\mathbb{F}_{p}$, where $S$ is an isotropic subspace of $\mathbb{F}_{p}^{2n}$.
Next, let us treat homogeneous subgroups of $\mathbb{H}^{n}(\mathbb{F}_{q})$. If $G$ is a homogeneous subgroup of $\mathbb{H}^{n},$ then $\pi_{h}(G)$ is a subspace of $\mathbb{F}_{q}^{2n}$ (over $\mathbb{F}_{q}$). So fix a subspace $S$ of $\mathbb{F}_{q}^{2n}$ and find all subgroups $G$ of $\mathbb{H}^{n}(\mathbb{F}_{q})$ such that $\pi_{h}(G)=S.$ Assume that $q=p^{r}$ for some prime $p$ and positive integer $r.$ The projection $\pi_{h}:\mathbb{H}^{n}\rightarrow\mathbb{F}_{q}^{2n}$ induces a group homomorphism $\pi_{h}|_{G}:$ $G\rightarrow S.$ Then elements of the kernel $K$ of $\pi_{h}|_{G}$ are of the form $(0,t)$ for $t\in\mathbb{F}_{q}$. Also, note that the restriction $\pi_{v}|_{K}:K\rightarrow\mathbb{F}_{q}$ embeds $K$ into $\mathbb{F}_{q}$ as a subgroup. Thus $|K|$ divides $q=p^{r}$, and so $|K|=p^{\ell}$ for some $\ell\leq r.$ Firstly, assume that $K$ is nontrivial, i.e., contains an element $(0,t)$ with $t\neq 0.$ Since $K$ is homogeneous, $K$ includes $\mathcal{O}_{\mathbb{F}_{q}^{*}}(0,t)$. So, by Lemma 2.3 (ii), we obtain $|K|\geq\frac{q-1}{2}=\frac{p^{r}-1}{2}$. This forces $|K|=p^{r}=q$ since $p$ is odd. (Note that if $\ell\leq r-1$, then it would be that $|K|<\frac{q-1}{2}.$) Since $|G|=|S||K|=|S|q,$ in this case we have $G=S\times\mathbb{F}_{q}.$ Secondly, assume that $K=\{(0,0)\}$ is trivial. Then $\pi_{h}|_{G}:G\rightarrow S$ is an isomorphism, and so, as in Proposition 2.4, $S$ is isotropic and $G=\Gamma_{S}(\rho)$ for some group homomorphism(equivalently, a linear map over $\mathbb{F}_{p}$) $\rho:S\rightarrow\mathbb{F}_{q}.$ We claim that $\rho$ is trivial. To show this, let $(x,\rho(x))\in\Gamma_{S}(\rho)=G$ with $0\neq x\in\mathbb{F}_{q}^{2n}$. Then by the homogeneity condition, for any $s\in\mathbb{F}_{q}^{*}$, we have
$s\cdot(x,\rho(x))=(sx,s^{2}\rho(x))=(sx,\rho(sx)).$ (16)
But, in particular, if $s\in\mathbb{F}_{p}^{*}$, then $\rho(sx)=s\rho(x)$ since $\rho$ is linear over $\mathbb{F}_{p}.$ By this fact and (16), we obtain the equality $s^{2}\rho(x)=s\rho(x)$ for any $s\in\mathbb{F}_{p}^{*}$. This implies that $\rho(x)=0$ for all nonzero $x\in\mathbb{F}_{q}^{2n}$, i.e., $\rho$ is trivial. Therefore, $G=\Gamma_{S}(\rho)=S\times\{0\}.$ In conclusion, we have proved:
Proposition 2.5.
If $G$ is a homogeneous subgroup of $\mathbb{H}^{n}(\mathbb{F}_{q})$, then $G$ is one of the following two cases
  • (a)
    $G=S\times\mathbb{F}_{q}$ for a subspace $S$ of $\mathbb{F}_{q}^{2n}$.
  • (b)
    $G=S\times\{0\}$ for an isotropic subspace $S$ of $\mathbb{F}_{q}^{2n}$.
2.2 Orthogonal complement of a homogeneous subgroup of $\mathbb{H}^{n}(\mathbb{F}_{q})$
For a subspace $V$ of the vector space $\mathbb{F}_{q}^{m},$ let $V^{\perp}$ be the subspace of $V$ defined as $V^{\perp}=\{w\in\mathbb{F}_{q}^{m}:\langle v,w\rangle=0\,\,\mathrm{for\,\,\,all}\,v\in V\}$, where $\langle\,$$\rangle$ is the standard inner product on $\mathbb{F}_{q}^{m}.$ We may call $V^{\perp}$ the orthogonal complement of $V$ w.r.t. $\langle\,$$\rangle$. By Proposition 2.5, homogeneous subgroups $G$ of $\mathbb{H}^{n}(\mathbb{F}_{q})$ are some subspaces of $\mathbb{F}_{q}^{2n+1},$ and so we can take $G^{\perp}$ for homogeneous subgroups $G$ of $\mathbb{H}^{n}(\mathbb{F}_{q})$. If $G=S\times\{0\}$ for an isotropic $S$ is a homogeneous subgroup, then we have $G^{\perp}=S^{\perp}\times\mathbb{F}_{q}$, which is necessarily a (vertical) homogeneous subgroup of $\mathbb{H}^{n}(\mathbb{F}_{q})$. If $G=S\times\mathbb{F}_{q}$, where $S$ is an arbitrary subspace of $\mathbb{F}_{q}^{2n}$, then $G^{\perp}=S^{\perp}\times\{0\}$. In this case $G^{\perp}$ is not necessarily a homogeneous subgroup of $\mathbb{H}^{n}(\mathbb{F}_{q})$. For example, if $\mathrm{dim}\,S\leq n-1$, then $\mathrm{dim}\,S^{\perp}\geq n+1$ and so $S$ is not isotropic by the dimension reason. Thus $G^{\perp}=S^{\perp}\times\{0\}$ is not a homogeneous subgroup of $\mathbb{H}^{n}(\mathbb{F}_{q})$. (In fact, it is not even a subgroup of $\mathbb{H}^{n}(\mathbb{F}_{q})$.)
2.3 Counting subgroups of $\mathbb{H}^{n}(\mathbb{F}_{q})$
Now we count subgroups of $\mathbb{H}^{n}(\mathbb{F}_{q})$ for the two special cases in the above. For $k=0,\ldots,2n$, let $\mathrm{Gr}(k,2n)$ be the variety of $k$ dimensional subspaces in $\mathbb{F}_{q}^{2n}.$ For $k=0,\ldots n,$ let $\mathrm{IG}(k,2n)$ be the variety of $k$ dimensional isotropic subspaces in $\mathbb{F}_{q}^{2n}$.
First we count subgroups of $\mathbb{H}^{n}(\mathbb{F}_{p})$ for a prime $p$, using Proposition 2.4. For a subspace $S$ of $\mathbb{F}_{p}^{2n}$ of dimension $k$ ($|S|=p^{k}$), let n(S)n(S)(resp. $n(S)^{\prime}$) be the number of subgroups $G$ of $\mathbb{H}^{n}(\mathbb{F}_{q})$ such that $\pi_{h}(G)=S$ and $|G|=p^{k}$ (resp. $|G|=p^{k+1}$). Then, by Proposition 2.4, the number of all subgroups of $\mathbb{H}^{n}(\mathbb{F}_{p})$ equals
$\sum_{k=0}^{2n}\sum_{S\in\mathrm{Gr}(k,2n)}|n(S)^{\prime}|+\sum_{k=0}^{n}\sum_{S\in\mathrm{IG}(k,2n)}|n(S)|.$
We easily see that $n(S)^{\prime}=1$ for each $S\in\mathrm{Gr}(k,2n)$; and $n(S)=kp$ for each $S\in\mathrm{IG}(k,2n)$ because there are $kp$ group homomorphisms (linear maps) from $S$ to $\mathbb{F}_{p}$, identified with $1\times k$ matrices with entries in $\mathbb{F}_{p}$. This number equals
$\sum_{k=0}^{2n}|\mathrm{Gr}(k,2n)|+\sum_{k=0}^{n}kp|\mathrm{IG}(k,2n)|.$ (17)
Next, we count homogeneous subgroups of $\mathbb{H}^{n}(\mathbb{F}_{q})$, using Proposition 2.5. For each subspaces $S$ of $\mathbb{F}_{p}^{2n}$ of dimension $k$ ($|S|=p^{k}$), let m(S)m(S)(resp. $m(S)^{\prime}$) be the number of homogeneous subgroups $G$ such that $\pi_{h}(G)=S$ and $|G|=p^{k}$ (resp. $|G|=p^{k+1}$). Then by Proposition 2.5, the number of all homogeneous subgroups of $\mathbb{H}^{n}(\mathbb{F}_{q})$ equals
$\sum_{k=0}^{2n}\sum_{S\in\mathrm{Gr}(k,2n)}|m(S)^{\prime}|+\sum_{k=0}^{n}\sum_{S\in\mathrm{IG}(k,2n)}|m(S)|.$
Since $m(S)^{\prime}=1$ for each $S\in\mathrm{Gr}(k,2n)$; and $m(S)=1$ for each $S\in\mathrm{IG}(k,2n)$, the number of all homogeneous subgroups of $\mathbb{H}^{n}(\mathbb{F}_{q})$ equals
$\sum_{k=0}^{2n}|\mathrm{Gr}(k,2n)|+\sum_{k=0}^{n}|\mathrm{IG}(k,2n)|.$ (18)
Now, by (17) and (18), enumerating subgroups of $\mathbb{H}^{n}(\mathbb{F}_{q})$ is reduced to counting elements of $|\mathrm{Gr}(k,2n)|$ and $|\mathrm{IG}(k,2n)|$. Let us give counting formulas for elements of these varieties. For an indeterminate $q$ and an integer $m>0$, let $m]_{q}$ be a polynomial in $q$
$m]_{q}:=\sum_{i=0}^{m-1}q^{i}=\frac{1-q^{m}}{1-q},$
and set
$m]_{q}!:=[m]_{q}[m-1]_{q}\cdots[2]_{q}[1]_{q}.$
Then the following formulas can be found in the literature, for example, in [22, 5] .
Proposition 2.6.
The cardinalities of $\mathrm{Gr}(k,2n)$ and $\mathrm{IG}(k,2n)$, respectively, are given by
$|\mathrm{Gr}(k,2n)|=\frac{[2n]_{q}!}{[k]_{q}![(2n-k)]_{q}!},$
$|\mathrm{IG}(k,2n)|=\frac{[n]_{q}!}{[k]_{q}![n-k]_{q}!}\cdot\prod_{i=n-k+1}^{n}\left(q^{i}+1\right).$
In Proposition 2.5, we dealt with only homogeneous subgroups of $\mathbb{H}^{n}(\mathbb{F}_{q})$. Using the idea in Propositions 2.4 and 2.5 we can still construct some subgroups of $\mathbb{H}^{n}(\mathbb{F}_{q})$. For example, fix a subgroup $S$ of $\mathbb{F}_{q}^{2n}$ so that $S$ is a subspace over $\mathbb{F}_{p}$. Let $F(S)$ be the smallest subfield of $\mathbb{F}_{q}$ containing all components $x_{i}\in\mathbb{F}_{q}$ of $x=(x_{1},\ldots,x_{2n})$, where $x$ runs through $S$. Then for any subfield $F$ of $\mathbb{F}_{q}$ with $F(S)\subset F$, the product $S\times F$ is a subgroup of $\mathbb{H}^{n}(\mathbb{F}_{q})$. Also, if $S$ is isotropic, then as in Proposition 2.4, we can obtain a subgroup $G$ of $\mathbb{H}^{n}(\mathbb{F}_{q})$ as a graph of a group homomorphism. The group thus constructed is of the non-product type. In case $S$ is not isotropic, we could not find a subgroup $G$ of the non-product type, with $\pi_{h}(G)=S$, and we do not know whether there exists such a subgroup $G$. It would be interesting to investigate existence of such subgroups of the non-product type, with $\pi_{h}(G)=S$ for a non-isotropic subgroup $S$, or, more generally, to classify and count all subgroups of $\mathbb{H}^{n}(\mathbb{F}_{q})$.
3 Proof of Theorem 1.2
Define a bilinear operator (form) $\mathcal{L}$ on $\mathscr{F}(\mathbb{F}_{q}^{2})$ by
$\mathcal{L}(f_{1},f_{2}):=\frac{1}{q^{3}}\sum_{x=(x_{1},x_{2}),y=(y_{1},y_{2})\in\mathbb{F}_{q}^{2}}1_{x_{1}\cdot y_{1}+y_{2}=x_{2}}f_{1}(x)f_{2}(y),$ (19)
for any $f_{1},f_{2}\in\mathscr{F}(\mathbb{F}_{q}^{2}).$ Let $\langle\,$$\rangle$ be the ‘standard’ inner product on $\mathscr{F}(\mathbb{F}_{q}^{2})$, i.e.,
$\langle f_{1},f_{2}\rangle=\frac{1}{q^{2}}\sum_{x=(x_{1},x_{2})\in\mathbb{F}_{q}^{2}}f_{1}(x)f_{2}(x).$
Then we can write $\mathcal{L}(f_{1},f_{2})=\langle f_{1},Af_{2}\rangle$, where $A$ is an endomorphism on $\mathscr{F}(\mathbb{F}_{q}^{2})$ defined by
$\left(Af_{2}\right)(x):=\frac{1}{q}\sum_{y=(y_{1},y_{2})\in\mathbb{F}_{q}^{2}}1_{x_{1}y_{1}+y_{2}=x_{2}}f_{2}(y)$
for $f\in\mathscr{F}(\mathbb{F}_{q}^{2})$. Note that the bilinear operator $\mathcal{L}$ is not symmetric.
For $1\leq r,s\leq\infty,$ denote by $A(s\to r)$ the smallest number, if it exist, such that the inequality
$\|Af\|_{L^{r}(\mathbb{F}^{2}_{q},dx)}\leq A(s\to r)\|f\|_{L^{s}(\mathbb{F}^{2}_{q},dx)}$
holds for all nonnegative $f\in\mathscr{F}(\mathbb{F}_{q}^{2}).$
To prove Theorem 1.2, we introduce a constant $L(u_{1},u_{2})$ as an alternative to the constant $LW(u_{1},u_{2}).$
Definition 3.1.
For $1\leq u_{1},u_{2}\leq\infty,$ define $L(u_{1},u_{2})$ as the smallest positive constant, if it exist, such that the following estimate holds for all nonnegative $f_{1},f_{2}\in\mathscr{F}(\mathbb{F}_{q}^{2})$
$\mathcal{L}(f_{1},f_{2})\leq L(u_{1},u_{2})\|f_{1}\|_{L^{u_{1}}(\mathbb{F}^{2}_{q},dx)}\cdot\|f_{2}\|_{L^{u_{2}}(\mathbb{F}^{2}_{q},dx)},$ (20)
and if it does not exist, we set $L(u_{1},u_{2}):=\infty.$
Recall that for $1\leq u\leq\infty$, the Holder’s conjugate of $u$ is defined by $\frac{1}{u}+\frac{1}{u^{\prime}}=1$.
Lemma 3.2.
If $1\leq u_{1},u_{2}\leq\infty$ satisfy $A(u_{2}\to u_{1}^{\prime})\lesssim 1$, then we have $L(u_{1},u_{2})\lesssim 1.$
Proof.
By Hölder’s inequality and the assumption, we have
$\mathcal{L}(f_{1},f_{2})=\langle f_{1},Af_{2}\rangle\leq\|f_{1}\|_{L^{u_{1}}(\mathbb{F}^{2}_{q},dx)}\,\|Af_{2}\|_{L^{u^{\prime}_{1}}(\mathbb{F}^{2}_{q},dx)}\lesssim\|f_{1}\|_{L^{u_{1}}(\mathbb{F}^{2}_{q},dx)}\,\|f_{2}\|_{L^{u_{2}}(\mathbb{F}^{2}_{q},dx)}.$
Since these inequalities holds for any $f_{1},f_{2}\in\mathscr{F}(\mathbb{F}_{q}^{2})$, the lemma is immediate. ∎
Theorem 3.3.
For $1\leq u_{1},u_{2}\leq\infty,$ the following holds.
$L(u_{1},u_{2})\lesssim 1\quad\mbox{if and only if}\quad\frac{1}{u_{1}}+\frac{2}{u_{2}}\leq 2\quad\mathrm{and}\quad\frac{2}{u_{1}}+\frac{1}{u_{2}}\leq 2.$
Proof.
First, assume $L(u_{1},u_{2})\lesssim 1$. To obtain the first inequality $\tfrac{1}{u_{1}}+\tfrac{2}{u_{2}}\leq 2$, take two test functions
$f_{1}(x_{1},x_{2})=\mathbf{1}_{\{x_{2}=x_{1}+1\}},\quad f_{2}(x_{1},x_{2})=\mathbf{1}_{\{x_{1}=x_{2}=1\}}.$
By simple calculations, we obtain
$\mathcal{L}(f_{1},f_{2})=q^{-2};$
$\|f_{1}\|_{L^{u_{1}}(\mathbb{F}^{2}_{q},dx)}=\Bigl(\frac{q}{q^{2}}\Bigr)^{1/u_{1}}=q^{-1/u_{1}},\qquad\|f_{2}\|_{L^{u_{2}}(\mathbb{F}^{2}_{q},dx)}=q^{-2/u_{2}}.$
The assumption $L(u_{1},u_{2})\lesssim 1$ gives
$q^{-2}\lesssim q^{-1/u_{1}-2/u_{2}},$
which is equivalent to the inequality $\tfrac{1}{u_{1}}+\tfrac{2}{u_{2}}\leq 2$.
The second inequality $\tfrac{2}{u_{1}}+\tfrac{1}{u_{2}}\leq 2$ can be obtained in a similar manner by choosing
$g_{1}(x_{1},x_{2})=\mathbf{1}_{\{x_{1}=x_{2}=1\}},\quad g_{2}(x_{1},x_{2})=\mathbf{1}_{\{x_{2}=1-x_{1}\}}.$
For the other direction, assume that $1\leq u_{1},u_{2}\leq\infty$ satisfy the inequalities
$\frac{1}{u_{1}}+\frac{2}{u_{2}}\leq 2\quad\mathrm{and}\quad\frac{2}{u_{1}}+\frac{1}{u_{2}}\leq 2.$
Let us show that $L(u_{1},u_{2})\lesssim 1$. By the interpolation theorem and the nesting property of the norm, it suffices to establish the estimates on the critical end-points $(\frac{1}{u_{1}},\frac{1}{u_{2}})\in[0,1]\times[0,1]$, which are $(0,1),(1,0),$ and $(\frac{2}{3},\frac{2}{3})$. In other words, it remains to prove the following estimates:
$L(\infty,1)\lesssim 1,~L(1,\infty)\lesssim 1,~L\left(\frac{3}{2},~\frac{3}{2}\right)\lesssim 1.$
But these estimates follows directly from Lemma 3.2 and the following result, due to Koh [16, Theorem 1.1]:
$A(1\to 1)\lesssim 1,~~A(\infty\to\infty)\lesssim 1,~~A\left(\frac{3}{2}\to 3\right)\lesssim 1.$
The following result is a special case of Theorem 3.3, but it is very useful in practice.
Corollary 3.4.
We have $L\left(\frac{3}{2},\frac{3}{2}\right)\lesssim 1.$
Proof.
The proof follows immediately from Theorem 3.3 since
$\frac{1}{u_{1}}+\frac{2}{u_{2}}=\frac{2}{u_{1}}+\frac{1}{u_{2}}=2$
with $u_{1}=u_{2}=\frac{3}{2}$. ∎
Now we can compare two constants $LW(u_{1},u_{2})$ and $L(u_{1},u_{2}).$
Lemma 3.5.
For $1\leq u_{1},u_{2}\leq\infty$, the following holds.
$L(u_{1},u_{2})\lesssim 1\quad\mathrm{if\,\,and\,\,only\,\,if}\quad LW(u_{1},u_{2})\lesssim 1.$
Proof.
To prove the lemma, it is sufficient to show the LHS of the inequality (7) with $n=1$ is equal to LHS, $\mathcal{L}(f_{1},f_{2})$, of the inequality (20). In fact, it follows from (1) that
$\frac{1}{q^{3}}\sum_{(x_{1},x_{2},t)\in\mathbb{H}_{q}^{1}}f_{1}(\pi_{1}(x_{1},x_{2},t))f_{2}(\pi_{2}(x_{1},x_{2},t))=\frac{1}{q^{3}}\sum_{x_{1},x_{2},t\in\mathbb{F}_{q}}f_{1}(x_{2},t+x_{1}x_{2})f_{2}(x_{1},t)=\mathcal{L}(f_{1},f_{2}),$
as desired. ∎
In conclusion, Theorem 1.2 follows from Theorem 3.3 and Lemma 3.5
4 Proof of Theorem 1.4
While proving Theorem 1.4, we will heavily use the following result.
Theorem 4.1.
Fix any $k=1,\ldots,n$. Then for any nonnegative $f_{1},\ldots,f_{2n}\in\mathscr{F}(\mathbb{F}_{q}^{2n}),$ the inequality
$\frac{1}{q^{2n+1}}\sum_{(x,t)\in\mathbb{H}^{n}(\mathbb{F}_{q})}\prod_{j=1}^{2n}f_{j}(\pi_{j}(x,t))$
$\lesssim\|f_{k}\|_{L^{\frac{2n+1}{2}}(\mathbb{F}_{q}^{2n},dx)}\|f_{n+k}\|_{L^{\frac{2n+1}{2}}(\mathbb{F}_{q}^{2n},dx)}\prod_{j=1,j\neq k}^{n}\left(\|f_{j}\|_{L^{2n+1}(\mathbb{F}_{q}^{2n},dx)}\|f_{n+j}\|_{L^{2n+1}(\mathbb{F}_{q}^{2n},dx)}\right)$ (21)
hold.
Proof.
We may assume, WLOG, that $k=1$. Denote
$I:=\frac{1}{q^{2n+1}}\sum_{(x,t)\in\mathbb{H}^{n}(\mathbb{F}_{q})}\prod_{j=1}^{2n}f_{j}(\pi_{j}(x,t)).$
Step $1^{\circ}$ (Separate the factor $f_{n}$.) For convenience, for $x=(x_{1},\ldots,x_{2n})\in\mathbb{F}_{q}^{2n}$, write $x=(x^{\prime},y)$, where $x^{\prime}=(x_{1},\ldots,x_{2n-1})\in\mathbb{F}_{q}^{2n-1}$ and $y=x_{2n}\in\mathbb{F}_{q}$, and the change of variables $\tau:=t-\frac{1}{2}x_{n}y$. Then we can write
$I$ $=\frac{1}{q^{2n+1}}\sum_{(x^{\prime},y,t)\in\mathbb{H}_{q}^{n}}f_{2n}(x^{\prime},t-\frac{1}{2}x_{n}y)f_{n}(\widehat{x}_{n},t+\frac{1}{2}x_{n}y)\,\prod_{\begin{subarray}{c}j=1\\ j\neq n,2n\end{subarray}}^{2n}f_{j}\big(\pi_{j}(x^{\prime},y,t)\big)$
$=\frac{1}{q^{2n+1}}\sum_{(x^{\prime},y,\tau)\in\mathbb{F}_{q}^{2n+1}}f_{2n}(x^{\prime},\tau)\,f_{n}(\widehat{x}_{n},\tau+x_{n}y)\prod_{\begin{subarray}{c}j=1\\ j\neq n,2n\end{subarray}}^{2n}f_{j}\big(\pi_{j}(x^{\prime},y,\tau+\frac{1}{2}x_{n}y)\big)$
$=\frac{1}{q^{2n}}\sum_{(x^{\prime},\tau)\in\mathbb{F}_{q}^{2n}}f_{2n}(x^{\prime},\tau)\,\frac{1}{q}\sum_{y\in\mathbb{F}_{q}}f_{n}(\widehat{x}_{n},\tau+x_{n}y)\prod_{\begin{subarray}{c}j=1\\ j\neq n,2n\end{subarray}}^{2n}f_{j}\big(\pi_{j}(x^{\prime},y,\tau+\frac{1}{2}x_{n}y)\big).$ (22)
Note that $f_{2n}$ is independent of $y$. By applying Hölder’s inequality with the conjugate pair of exponents $(2n+1,\frac{2n+1}{2n})$, we obtain
$I\leq$ $\bigg(\frac{1}{q^{2n}}\sum_{(x^{\prime},\tau)\in\mathbb{F}_{q}^{2n}}f_{2n}(x^{\prime},\tau)^{2n+1}\bigg)^{\frac{1}{2n+1}}\cdot J=\|f_{2n}\|_{L^{2n+1}(\mathbb{F}_{q}^{2n},dx)}\cdot J,$
where $J$ captures the contribution of the remaining product
$J:=\left[\frac{1}{q^{2n}}\sum_{(x^{\prime},\tau)\in\mathbb{F}_{q}^{2n}}\bigg(\frac{1}{q}\sum_{y\in\mathbb{F}_{q}}f_{n}(\widehat{x}_{n},\tau+x_{n}y)\prod_{\begin{subarray}{c}j=1\\ j\neq n,2n\end{subarray}}^{2n}f_{j}\big(\pi_{j}(x^{\prime},y,\tau+\frac{1}{2}x_{n}y)\big)\bigg)^{\frac{2n+1}{2n}}\right]^{\frac{2n}{2n+1}}.$
To prove (4.1), it is enough to show that
$J\lesssim\|f_{1}\|_{L^{\frac{2n+1}{2}}(\mathbb{F}_{q}^{2n},dx)}\|f_{n+1}\|_{L^{\frac{2n+1}{2}}(\mathbb{F}_{q}^{2n},dx)}\|f_{n}\|_{L^{2n+1}(\mathbb{F}_{q}^{2n},dx)}\prod_{j=2}^{n-1}\left(\|f_{j}\|_{L^{2n+1}(\mathbb{F}_{q}^{2n},dx)}\|f_{n+j}\|_{L^{2n+1}(\mathbb{F}_{q}^{2n},dx)}\right).$
Step $2^{\circ}$ (Extract the $f_{n}$-factor). For fixed $y$ and $x_{n}$, perform the change of variables $s:=\tau+x_{n}y$. For $x=(x_{1},\ldots,x_{2n})\in\mathbb{F}_{q}^{2n}$, we use the notation $\widehat{x}_{i,j}$ to denote the element in $\mathbb{F}_{q}^{2n-2}$ obtained by removing $x_{i}$ and $x_{j}$ from $x$. By Minkowski inequality, we obtain
$J=$ $\left[\frac{1}{q^{2n-1}}\sum_{x^{\prime}\in\mathbb{F}_{q}^{2n-1}}\frac{1}{q}\sum_{s\in\mathbb{F}_{q}}\bigg(\frac{1}{q}\sum_{y\in\mathbb{F}_{q}}f_{n}(\widehat{x}_{n},s)\prod_{\begin{subarray}{c}j=1\\ j\neq n,2n\end{subarray}}^{2n}f_{j}\big(\pi_{j}(x^{\prime},y,s-\frac{1}{2}x_{n}y)\big)\bigg)^{\frac{2n+1}{2n}}\right]^{\frac{2n}{2n+1}}$
$\leq$ $\frac{1}{q}\sum_{y\in\mathbb{F}_{q}}\bigg[\frac{1}{q^{2n-1}}\sum_{x^{\prime}\in\mathbb{F}^{2n-1}_{q}}\frac{1}{q}\sum_{s\in\mathbb{F}_{q}}f_{n}(\widehat{x}_{n},s)^{\frac{2n+1}{2n}}\prod_{\begin{subarray}{c}j\neq n,2n\end{subarray}}f_{j}\big(\pi_{j}(x^{\prime},y,s-\frac{1}{2}x_{n}y)\big)^{\frac{2n+1}{2n}}\bigg]^{\frac{2n}{2n+1}}$
$=$ $\frac{1}{q}\sum_{y\in\mathbb{F}_{q}}\bigg[\frac{1}{q^{2n-1}}\sum_{(\widehat{x}_{n,2n},s)\in\mathbb{F}^{2n-1}_{q}}f_{n}(\widehat{x}_{n},s)^{\frac{2n+1}{2n}}\bigg(\frac{1}{q}\sum_{x_{n}\in\mathbb{F}_{q}}\prod_{\begin{subarray}{c}j\neq n,2n\end{subarray}}f_{j}\big(\pi_{j}(x^{\prime},y,s-\frac{1}{2}x_{n}y)\big)^{\frac{2n+1}{2n}}\bigg)\bigg]^{\frac{2n}{2n+1}}$
where the last equality we have used the fact that $f_{n}$-term is independent of $x_{n}$. By applying Hölder inequality with the conjugate pair of exponents $(2n,\frac{2n}{2n-1})$, we obtain
$J\leq\frac{1}{q}\sum_{y\in\mathbb{F}_{q}}J_{1}(y)J_{2}(y)$
where
$J_{1}(y)=$ $\bigg(\frac{1}{q^{2n-1}}\sum_{(\widehat{x}_{n,2n},s)\in\mathbb{F}^{2n-1}_{q}}f_{n}(\widehat{x}_{n},s)^{2n+1}\bigg)^{\frac{1}{2n+1}},$
$J_{2}(y)=$ $\bigg(\frac{1}{q^{2n-1}}\sum_{(\widehat{x}_{n,2n},s)\in\mathbb{F}_{q}^{2n-1}}\bigg(\frac{1}{q}\sum_{x_{n}\in\mathbb{F}_{q}}\prod_{\begin{subarray}{c}j\neq n,2n\end{subarray}}f_{j}\big(\pi_{j}(x^{\prime},y,s-\frac{1}{2}x_{n}y)^{\frac{2n+1}{2n}}\bigg)^{\frac{2n}{2n-1}}\bigg)^{\frac{2n-1}{2n+1}}$
Applying Hölder inequality in the outer yy–sum with the conjugate pair of exponents $\left(2n+1,\frac{2n+1}{2n}\right)$ yields
$J\lesssim$ $\Big(\frac{1}{q}\sum_{y}J_{1}(y)^{2n+1}\Big)^{\frac{1}{2n+1}}\cdot\Bigg(\frac{1}{q}\sum_{y\in\mathbb{F}_{q}}J_{2}(y)^{\frac{2n+1}{2n}}\Bigg)^{\frac{2n}{2n+1}}$
$=$ $\bigg(\frac{1}{q^{2n}}\sum_{(\widehat{x}_{n},s)\in\mathbb{F}^{2n}_{q}}f_{n}(\widehat{x}_{n},s)^{2n+1}\bigg)^{\frac{1}{2n+1}}\cdot J_{\Pi}$
$=$ $\|f_{n}\|_{L^{2n+1}(\mathbb{F}_{q}^{2n},dx)}\cdot J_{\Pi}.$
Therefore it remains to bound
$J_{\Pi}:=\Bigg(\frac{1}{q}\sum_{y\in\mathbb{F}_{q}}J_{2}(y)^{\frac{2n+1}{2n}}\Bigg)^{\frac{2n}{2n+1}}$
by the product of the remaining norms appearing on the right-hand side of (4.1).
Step $3^{\circ}$ (Apply the induction hypothesis).
By applying Minkowski’s inequality, we obtain
$J_{\Pi}=$ $\Bigg(\frac{1}{q}\sum_{y\in\mathbb{F}_{q}}J_{2}(y)^{\frac{2n+1}{2n}}\Bigg)^{\frac{2n}{2n+1}}$
$\lesssim$ $\left(\frac{1}{q^{2}}\sum_{x_{n},y\in\mathbb{F}_{q}}\left[\frac{1}{q^{2n-1}}\sum_{(\widehat{x}_{n,2n},t)\in\mathbb{F}_{q}^{2n-1}}\prod_{\begin{subarray}{c}j\neq n,2n\end{subarray}}f_{j}(\pi_{j}(x^{\prime},y,t))^{\frac{2n+1}{2n-1}}\right]^{\frac{2n-1}{2n}}\right)^{\frac{2n}{2n+1}}.$ (23)
Observing that the expression inside the square brackets of (4) is a product of 2n22n-2 functions, we aim to transform it into a form suitable for applying the induction hypothesis to establish an upper bound for $J_{\Pi}$. We recall that
$f_{j}(\pi_{j}(x^{\prime},y,\tau))=\begin{cases}f_{j}(\check{x}_{j},\tau+\frac{1}{2}x_{j}x_{j+n}),~~~\text{ if }~~~~~~1\leq j\leq n-1,\\ f_{j}(\check{x}_{n+j},\tau-\frac{1}{2}x_{j-n}x_{j}),~~~\text{ if }n+1\leq j\leq 2n-1.\end{cases}$ (24)
We temporarily represent points in $\mathbb{H}^{n-1}(\mathbb{F}_{q})$ using coordinates $(u,t)=(u_{1},\ldots,u_{2n-2},t)$ where $u\in\mathbb{F}_{q}^{2n-2}$. Fixing $x_{n},y\in\mathbb{F}_{q}$ we define the functions $g_{j}$, $j\in\{1,\ldots,2n-2\}$, on $\mathbb{F}_{q}^{2n-2}$ by
$g_{j}(\widehat{u}_{j},t)=\begin{cases}f_{j}(\tilde{u}_{j})^{\tfrac{2n+1}{2n-1}},~~~~~~~~~~1\leq j\leq n-1,\\ f_{j+1}(\tilde{u}_{j})^{\tfrac{2n+1}{2n-1}},~~~~~n\leq j\leq 2n-2,\end{cases}$
where $\tilde{u}_{j}$ is obtained from
$(u_{1},u_{2},\ldots,u_{n-1},x_{n},u_{n},\ldots,u_{2n-2},y,t)$
by removing $u_{j}$. More precisely
$g_{j}(\widehat{u}_{j},t)=\begin{cases}f_{1}(u_{2},\ldots,u_{n-1},x_{n},u_{n},\ldots,u_{2n-2},y,t)^{\frac{2n+1}{2n-1}},~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~j=1,\\ f_{j}(u_{1},\ldots,u_{j-1},u_{j+1},\ldots,u_{n-1},x_{n},u_{n},\ldots,u_{2n-2},y,t)^{\frac{2n+1}{2n-1}},~~~2\leq j\leq n-2,\\ f_{n-1}(u_{1},\ldots,u_{n-2},x_{n},u_{n},\ldots,u_{2n-2},y,t)^{\frac{2n+1}{2n-1}},~~~~~~~~~~~~~~~~~~~~~~~~~j=n-1,\end{cases}$ (25)
and
$g_{j}(\widehat{u}_{j},t)=\begin{cases}f_{n+1}(u_{1},\ldots,u_{n-1},x_{n},u_{n+1},\ldots,u_{2n-2},y,t)^{\frac{2n+1}{2n-1}},~~~~~~~~~~~~~~~~~~~~~~~~~j=n,\\ f_{j}(u_{1},\ldots,u_{n-1},x_{n},u_{n},\ldots,u_{j-1},u_{j+1},\ldots,u_{2n-2},y,t)^{\frac{2n+1}{2n-1}},~n+1\leq j\leq 2n-3,\\ f_{2n-1}(u_{1},\ldots,u_{n-2},x_{n},u_{n},\ldots,u_{2n-3},y,t)^{\frac{2n+1}{2n-1}},~~~~~~~~~~~~~~~~~~~~~~~~~j=2n-2.\end{cases}$ (26)
With this notation in place, and recalling (24), we can restate (4) equivalently as follows
$J_{\Pi}\leq$ $\left(\frac{1}{q^{2}}\sum_{x_{n},y\in\mathbb{F}_{q}}\left[\frac{1}{q^{2n-1}}\sum_{(u,t)\in\mathbb{F}_{q}^{2n-1}}\prod_{\begin{subarray}{c}1\leq j\leq 2n-2\end{subarray}}g_{j}(\pi_{j}(u,t))\right]^{\frac{2n-1}{2n}}\right)^{\frac{2n}{2n+1}}$
where we recall that $\pi_{j}$ is the projection from $\mathbb{H}^{n-1}(\mathbb{F}_{q})$ to the vertical plane $W_{j}$ (identified with $\mathbb{F}_{q}^{2n-2}$), defined in (5). By the induction hypothesis we have
$\frac{1}{q^{2n-1}}$ $\sum_{(u,t)\in\mathbb{F}_{q}^{2n-1}}\prod_{\begin{subarray}{c}1\leq j\leq 2n-2\end{subarray}}g_{j}(\pi_{j}(u,t))$
$\;\lesssim\;$ $\|g_{1}\|_{L^{\frac{2n-1}{2}}(\mathbb{F}_{q}^{2n-2},dx)}\,\|g_{n}\|_{L^{\frac{2n-1}{2}}(\mathbb{F}_{q}^{2n-2},dx)}\,\prod_{\begin{subarray}{c}1\leq j\leq 2n-2,\\ j\neq 1,n\end{subarray}}\|g_{j}\|_{L^{2n-1}(\mathbb{F}_{q}^{2n-2},dx)}.$ (27)
Next we apply the multilinear Hölder’s inequality with exponents $\ell_{1}=\ell_{n}=n$ and $\ell_{j}=2n$, $\forall\,j\neq 1,n$. Note that $\sum\limits_{j=1}^{2n-2}\frac{1}{\ell_{j}}=1$. Hence, from (4), we have
$J_{\Pi}\lesssim$ $\bigg[\bigg(\frac{1}{q^{2}}\sum_{x_{n},y\in\mathbb{F}_{q}}\|g_{1}\|_{L^{\frac{2n-1}{2}}(\mathbb{F}_{q}^{2n-2},dx)}^{\frac{2n-1}{2}}\bigg)^{\frac{1}{n}}\cdot\bigg(\frac{1}{q^{2}}\sum_{x_{n},y\in\mathbb{F}_{q}}\|g_{n}\|_{L^{\frac{2n-1}{2}}(\mathbb{F}_{q}^{2n-2},dx)}^{\frac{2n-1}{2}}\bigg)^{\frac{1}{n}}$
$\times\prod_{\begin{subarray}{c}1\leq j\leq 2n-2,\\ j\neq 1,n\end{subarray}}\bigg(\frac{1}{q^{2}}\sum_{x_{n},y\in\mathbb{F}_{q}}\|g_{j}\|_{L^{2n-1}(\mathbb{F}_{q}^{2n-2},dx)}^{2n-1}\bigg)^{\frac{1}{2n}}\bigg]^{\frac{2n}{2n+1}}$
$=$ $\|f_{1}\|_{L^{\frac{2n+1}{2}}(\mathbb{F}_{q}^{2n},dx)}\|f_{n+1}\|_{L^{\frac{2n+1}{2}}(\mathbb{F}_{q}^{2n},dx)}\prod_{j=2}^{n-1}\left(\|f_{j}\|_{L^{2n+1}(\mathbb{F}_{q}^{2n},dx)}\|f_{n+j}\|_{L^{2n+1}(\mathbb{F}_{q}^{2n},dx)}\right).$
Here the last equality follows from the definition of $g_{j}$ for $j\in\{1,\ldots,2n-2\}$, given in (25) and (26).
For a positive $n$, we define a multilinear map
$T:\mathscr{F}(\mathbb{F}_{q}^{2n})\times\cdots\times\mathscr{F}(\mathbb{F}_{q}^{2n})\rightarrow\mathscr{F}(\mathbb{F}_{q}^{2n})$
as follows: For an element $(z,\tau)=(z_{1},\ldots,z_{2n-1},\tau)\in\mathbb{F}_{q}^{2n}$
$T(g_{1},\ldots,g_{2n-1})(z,\tau):=\frac{1}{q}\sum_{s\in\mathbb{F}_{q}}\left[g_{n}\big(z,\tau+z_{n}s\big)\cdot\prod_{\begin{subarray}{c}j=1,j\neq n\end{subarray}}^{2n-1}g_{j}\left(\pi_{j}\big(z,s,\tau+\frac{1}{2}z_{n}s\big)\right)\right].$
We use this operator $T$ to find an upper bound of $I$, given in (4). More precisely, taking a change of variable $\tau=t+\frac{1}{2}x_{n}x_{2n}$, and applying Hölder’s inequality with the conjugate pair of exponents
$(\ell,\ell^{\prime})=\bigg(\frac{n(2n+1)}{n+1},\frac{n(2n+1)}{2n^{2}-1}\bigg),$ (28)
we obtain
$I=$ $\frac{1}{q^{2n+1}}=\sum_{(x,t)\in\mathbb{H}^{n}(\mathbb{F}_{q})}\prod_{j=1}^{2n}f_{j}(\pi_{j}(x,t))$
$=$ $\frac{1}{q^{2n}}\sum_{(x^{\prime},t)\in\mathbb{F}_{q}^{2n}}\sum_{x_{2n}\in\mathbb{F}_{q}}f_{2n}(x^{\prime},t-\frac{1}{2}x_{n}x_{2n})\,\Bigg(\frac{1}{q}f_{n}(\widehat{x}_{n},t+\frac{1}{2}x_{n}x_{2n})\prod_{\begin{subarray}{c}j\neq n,2n\end{subarray}}f_{j}\big(\pi_{j}(x,t)\big)\Bigg)$
$=$ $\frac{1}{q^{2n}}\sum_{(x^{\prime},\tau)\in\mathbb{F}_{q}^{2n}}f_{2n}(x^{\prime},\tau)\,\Bigg(\frac{1}{q}\sum_{x_{2n}\in\mathbb{F}_{q}}f_{n}(\widehat{x}_{n},\tau-\frac{1}{2}x_{n}x_{2n})\prod_{\begin{subarray}{c}j\neq n,2n\end{subarray}}f_{j}\big(\pi_{j}(x^{\prime},x_{2n},\tau+\frac{1}{2}x_{n}x_{2n})\big)\Bigg)$
$\leq$ $\|T(f_{1},\ldots,f_{2n-1})\|_{L^{\ell^{\prime}}(\mathbb{F}^{2n}_{q},dx)}\cdot\|f_{2n}\|_{L^{\frac{n(2n+1)}{n+1}}(\mathbb{F}^{2n}_{q},dx)}.$ (29)
We are now ready to prove Theorem 1.4.
Proof of Theorem 1.4.
In view of (4), the proof of Theorem 1.4 is reduced to proving
$\|T(f_{1},\ldots,f_{2n-1})\|_{L^{\ell^{\prime}}(\mathbb{F}^{2n}_{q},dx)}\leq\prod_{j=1}^{2n-1}\|f_{j}\|_{L^{\ell_{j}}(\mathbb{F}^{2n}_{q},dx)}$ (30)
for all nonnegative $f_{1},\ldots,f_{2n-1}\in\mathscr{\mathbb{F}_{q}^{2n}}$, where $\ell^{\prime}$ is defined in (28) and $\ell_{1}=\ldots=\ell_{2n-1}=\frac{n(2n+1)}{n+1}$. To prove the inequality (30), first, from (4.1) we obtain
$\|T(f_{1},\ldots,f_{2n-1})\|_{L^{\ell_{k}}(\mathbb{F}^{2n}_{q},dx)}\lesssim\prod_{j=1}^{2n-1}\|f_{j}\|_{L^{\ell_{j,k}}(\mathbb{F}^{2n}_{q},dx)},~~~k=1,\ldots,n$ (31)
where $\ell_{k}$, $\ell_{j,k}$ are given by
$\ell_{k}=\begin{cases}\frac{2n+1}{2n},~~~~~~~~~~k=1,\ldots,n-1,\\ \frac{2n+1}{2n-1},~~~~~~~~~~k=n\end{cases}$ (32)
and
[…]
$j=\{1,\ldots,2n-1\}$, $k=1,\ldots,n$.
Note that the exponents in (32) and (33) satisfy
$\frac{1}{\ell^{\prime}}=\frac{2n^{2}-1}{n(2n+1)}=\frac{1}{n}\bigg[\frac{2n}{2n+1}\cdot(n-1)+\frac{2n-1}{2n}\bigg]=\sum\limits_{k=1}^{n}\frac{1}{n}\frac{1}{\ell_{k}}$
and
$\frac{1}{\ell_{j}}=\sum\limits_{k=1}^{n}\frac{1}{n}\frac{1}{\ell_{j,k}},~~~~~~~~j=1,\ldots,2n-1.$
To prove (30), we apply the discrete version of the multilinear interpolation theorem, a generalization to [17, Proposition 6.1]. Since $(\mathbb{F}_{q}^{2n},dx)$ is a finite measure space equipped with the normalized counting measure, the same interpolation argument carries over verbatim. We then perform (n1)(n-1) iterations with $\theta=\tfrac{1}{k}$, starting from $k=2$ and ending at $k=n$, using the identity
$\frac{1}{k}[a_{1}+\cdots+a_{k}]=\left(1-\frac{1}{k}\right)\left(\frac{1}{k-1}[a_{1}+\cdots+a_{k-1}]+\frac{1}{k}a_{k}\right).$
More precisely, we first apply multilinear interpolation theorem with $\theta=\frac{1}{2}$ to the two operator bounds given by (31) for $k=1$ and $k=2$. Next, we apply the multilinear interpolation theorem with $\theta=\frac{1}{3}$ to interpolate between the newly obtained bound and the operator bound stated in (31) for $k=3$. We continue this process until, in the final step, we apply the theorem with $\theta=\frac{1}{n}$ to interpolate between the previously obtained bound and the bound for $k=n$. From this, we obtain (30) for all nonnegative $f_{1},\ldots,f_{2n}\in\mathscr{F}(\mathbb{F}^{2n}_{q})$.
5 Proof of Corollary 1.5
Proof.
For subsets $K_{1},\ldots,K_{2n}\subset\mathbb{F}_{q}^{2n}$, denote
$K:=\bigcap\limits_{j=1}^{2n}\pi_{j}^{-1}(K_{j}),$
where each $K_{j}$ is viewed as a subset of $\mathbb{H}^{n}(\mathbb{F}_{q})$ through the identification $\mathbb{F}_{q}^{2n}=W_{j}\subset\mathbb{H}^{n}(\mathbb{F}_{q})$ for $j=1,\ldots 2n$. Then for $(x,t)\in\mathbb{H}^{n}(\mathbb{F}_{q}),$ we can write
$\mathbf{1}_{K}(x,t)=\prod_{j=1}^{2n}\mathbf{1}_{K_{j}}(\pi_{j}(x,t)).$
Then, to prove Corollary 1.5, it is sufficient to prove the inequality
$\frac{1}{q^{2n+1}}\sum_{(x,t)\in\mathbb{H}^{n}(\mathbb{F}_{q})}\prod_{j=1}^{2n}\mathbf{1}_{K_{j}}(\pi_{j}(x,t))\;\;\lesssim\;\;\prod_{j=1}^{2n}\|\mathbf{1}_{K_{j}}\|_{L^{\frac{n(2n+1)}{n+1}}(\mathbb{F}_{q}^{2n},dx)}.$ (34)
Here all $L^{p}$-norms are taken with respect to the normalized counting measure on $\mathbb{F}_{q}^{2n}$, so that
$\|\mathbf{1}_{K_{j}}\|_{L^{p}}=|K_{j}|^{1/p},\qquad j=1,\dots,2n.$
Finally, applying Theorem 1.4 with $f_{j}=\chi_{K_{j}}$ for $j=1,\ldots,2n$ immediately yields (34), as desired. ∎
6 Proof of Theorem 1.6
To prove Theorem 1.6, we invoke the following incidence bound, due to Vinh [23].
Theorem 6.1.
Let $\mathscr{P}$ be a set of points in $\mathbb{F}_{q}^{2}$ and $\mathscr{L}$ a set of lines in $\mathbb{F}_{q}^{2}$. The number of incidences between $\mathscr{P}$ and $\mathscr{L}$, denoted by $I(\mathscr{P},\mathscr{L})$, satisfies
$I(\mathscr{P},\mathscr{L})\leq\frac{|\mathscr{P}|\,|\mathscr{L}|}{q}+2q^{\frac{1}{2}}\sqrt{|\mathscr{P}|\,|\mathscr{L}|}.$
Proof of Theorem 1.6.
As in Section 3, we define a bijective map $\phi\colon\mathbb{F}_{q}^{3}\to\mathbb{F}_{q}^{3}$ by $\phi(x,y,t)=(x,y,t+\frac{xy}{2})$, then
$\frac{1}{q^{3}}\sum_{(x_{1},x_{2},t)\in\mathbb{H}_{q}^{1}}f_{1}(\pi_{1}(x_{1},x_{2},t))f_{2}(\pi_{2}(x_{1},x_{2},t))=\frac{1}{q^{3}}\sum_{x_{1},x_{2},t\in\mathbb{F}_{q}}f_{1}(\pi_{1}(\phi(x_{1},x_{2},t)))f_{2}(\pi_{2}(\phi(x_{1},x_{2},t)))$
$=\frac{1}{q^{3}}\sum_{x_{1},x_{2},t\in\mathbb{F}_{q}}f_{1}(x_{2},t+x_{1}x_{2})f_{2}(x_{1},t)$
$=\frac{1}{q^{3}}\sum_{(a_{1},a_{2}),(b_{1},b_{2})\in\mathbb{F}^{2}_{q},a_{1}b_{1}+b_{2}=a_{2}}f_{1}(a_{1},a_{2})f_{2}(b_{1},b_{2}).$
Let $f_{1}$ and $f_{2}$ be the indicator functions of $\pi_{1}(K)$ and $\pi_{2}(K)$, respectively. Note that the equation $a_{1}b_{1}+b_{2}=a_{2}$ is equivalent to an incidence between a point $(a_{1},a_{2})$ and a line defined by $y=b_{1}x+b_{2}$. Let $\mathscr{P}$ and $\mathscr{L}$ be the corresponding sets of points and lines, i.e. $|\mathscr{P}|=|\pi_{1}(K)|$ and $|\mathscr{L}|=|\pi_{2}(K)|$. Therefore, applying Theorem 6.1, one has
$\frac{1}{q^{3}}\sum_{(x_{1},x_{2},t)\in\mathbb{F}_{q}^{3}}f_{1}(\pi_{1}(x_{1},x_{2},t))f_{2}(\pi_{2}(x_{1},x_{2},t))=\frac{|K|}{q^{3}}\leq\frac{I(\pi_{1}(K),\pi_{2}(K))}{q^{3}},$
which is at most
$\frac{|\pi_{1}(K)||\pi_{2}(K)|}{q^{4}}+2q^{-\frac{5}{2}}\sqrt{|\pi_{1}(K)||\pi_{2}(K)|}.$
This implies the desired conclusion. ∎
7 Acknowledgements
Thang Pham and Dung The Tran would like to thank to the VIASM for the hospitality and for the excellent working condition. Daewoong Cheong sincerely thanks the Korea Institute for Advanced Study for financial support and hospitality and was partially supported by the research year program of Chungbuk National University in 2025.