### v2

# A functional Loomis–Whitney type inequality in the Heisenberg group and projection theorems over finite fields

Abstract
We establish functional Loomis–Whitney type inequalities in the finite Heisenberg group $\mathbb{H}^{n}(\mathbb{F}_{q})$. For $n=1$, we determine the sharp region of exponents $(u_{1},u_{2})$ for which the Heisenberg Loomis–Whitney inequality
$\frac{1}{q^{3}}\sum_{(x,t)\in\mathbb{H}^{1}(\mathbb{F}_{q})}f_{1}(\pi_{1}(x,t))\,f_{2}(\pi_{2}(x,t))\;\lesssim\;\|f_{1}\|_{L^{u_{1}}(\mathbb{F}_{q}^{2},dx)}\|f_{2}\|_{L^{u_{2}}(\mathbb{F}_{q}^{2},dx)}$
holds uniformly in $q$, namely
$\frac{1}{u_{1}}+\frac{2}{u_{2}}\leq 2\quad\text{and}\quad\frac{2}{u_{1}}+\frac{1}{u_{2}}\leq 2,$
which includes the endpoint estimate $L^{\frac{3}{2}}\times L^{\frac{3}{2}}\to L^{1}$. For general $n$, we prove the symmetric multilinear estimate at the endpoint exponent $u=\frac{n(2n+1)}{n+1},$ using an induction on $n$ that exploits the Heisenberg fiber structure together with a multilinear interpolation scheme. Specializing to indicator functions yields a sharp Loomis–Whitney type set inequality bounding $|K|$ for every finite $K\subset\mathbb{H}^{n}(\mathbb{F}_{q})$ in terms of the sizes of its $2n$ Heisenberg projections $\{\pi_{j}(K)\}_{j=1}^{2n}$, and in particular,
$\max_{1\leq j\leq 2n}|\pi_{j}(K)|\;\gtrsim_{n}\;|K|^{\frac{2n+1}{2(n+1)}}\,q^{-\frac{1}{2(n+1)}}.$
This result is optimal up to absolute constants. Moreover, when $n=1$ and $|K|>q$, we obtain a stronger statement via Vinh’s point–line incidence theorem. We also discuss connections to a boundedness problem for multilinear forms/operators over finite fields studied by Bhowmik, Iosevich, Koh, and Pham (2025), and to orthogonal projection/covering questions in $\mathbb{F}_{q}^{2n+1}$ studied by Chen (2018).
1 Introduction
The Loomis–Whitney inequality ([15]) bounds the volume of a dd–dimensional body in $\mathbb{R}^{d}$ by the volumes of its (d1)(d-1)–dimensional coordinate projections. In the discrete setting, it reads as follows. For $d\geq 2$, let $B_{1},\dots,B_{d}$ be finite sets, and let $A\subseteq B_{1}\times\cdots\times B_{d}$. For $i\in\{1,\ldots,d\}$, define
$\pi_{i}\colon A\to B_{1}\times\cdots\times B_{i-1}\times B_{i+1}\times\cdots\times B_{d},\qquad\pi_{i}(a_{1},\ldots,a_{d}):=(a_{1},\ldots,a_{i-1},a_{i+1},\ldots,a_{d}).$
Then
$|A|^{\,d-1}\leq\prod_{i=1}^{d}|\pi_{i}(A)|.$ (1)
There are powerful entropy-theoretic generalizations of (1), notably Han’s inequality and Shearer’s inequality [9, 16]. Variants of this inequality, in relation to set addition, also appear in additive combinatorics for abelian groups with numerous applications [8, 18].
This paper investigates Loomis–Whitney type inequalities in the noncommutative setting, focusing on the Heisenberg group over a finite field $\mathbb{F}_{q}$, where $q$ is an odd prime power. In this context, such inequalities are not only of independent interest but also shed light on group structures and are connected to structural properties of the Heisenberg group and to related combinatorial problems. Related works on the Heisenberg group can be found in [1, 11, 12, 17]. Interestingly, this topic is naturally connected to boundedness problems for multilinear forms/operators over finite fields and to orthogonal projection/covering questions in $\mathbb{F}_{q}^{2n+1}$. Further details will be discussed in Section 7.
In the continuous setting, the Loomis–Whitney type inequalities for the Heisenberg group (and, more generally, Carnot group) have been studied extensively with applications in multilinear restriction theory and related areas; for example, see [3, 7, 21] and references therein.
We now recall the definition of the Heisenberg group. For a positive integer $n$, the $n$-th Heisenberg group $\mathbb{H}^{n}(\mathbb{F}_{q})$ is the set $\mathbb{F}^{2n+1}_{q}$ equipped with the group multiplication defined as follows: for $(x,t),(x^{\prime},t^{\prime})\in\mathbb{F}^{2n+1}_{q}$ with $x,x^{\prime}\in\mathbb{F}_{q}^{2n}$ and $t,t^{\prime}\in\mathbb{F}_{q},$
$(x,t)\cdot(x^{\prime},t^{\prime})=\left(x+x^{\prime},t+t^{\prime}+\frac{1}{2}\sum\limits_{j=1}^{n}(x_{j}x^{\prime}_{n+j}-x_{n+j}x^{\prime}_{j})\right).$ (2)
By definition, the group $\mathbb{H}^{n}(\mathbb{F}_{q})$ and the vector space $\mathbb{F}_{q}^{2n+1}$ have the same underlying set. For each $1\leq j\leq 2n$, let
$W_{j}:=\{(x,t)\in\mathbb{F}_{q}^{2n+1}:x_{j}=0\},\qquad$ (3)
be the coordinate hyperplane corresponding to $x_{j}$, and let
$L_{j}:=\{(se_{j},0)\in\mathbb{F}_{q}^{2n+1}:s\in\mathbb{F}_{q}\}$
be the coordinate line corresponding to $x_{j}$, where $\{e_{j}\}_{j=1}^{2n}$ form the standard basis of $\mathbb{F}_{q}^{2n}$.
Then we easily see that $W_{j}$ and $L_{j}$ are subgroups of $\mathbb{H}^{n}(\mathbb{F}_{q})$ for $j\in\{1,\ldots,2n\}$. Further, $L_{j}$ is the orthogonal complement $W_{j}^{\perp}$ of $W_{j}$, with respect to the standard inner product on $\mathbb{F}_{q}^{2n+1}$; see Subsection 2.2 for these materials. The subgroups $W_{j}$ and $L_{j}$ will be often referred to as vertical and horizontal, respectively. We will provide more details on subgroup structures of $\mathbb{H}^{n}(\mathbb{F}_{q})$ in Section 2.
For $x\in\mathbb{F}^{2n}_{q}$ and $j\in\{1,\ldots,2n\}$, we use the notation $\widehat{x}_{j}$ to represent the element in $\mathbb{F}^{2n-1}_{q}$ obtained by entirely removing the $j$-th coordinate of $x$ while the notation $\check{x}_{j}$ represents the element in $\mathbb{F}^{2n}_{q}$ obtained by setting the $j$-th coordinate of $x$ to 0. For $j\in\{1,\ldots,n\}$, let
$\begin{cases}\pi_{j}:\mathbb{H}^{n}(\mathbb{F}_{q})\rightarrow W_{j}\\ \pi_{n+j}:\mathbb{H}^{n}(\mathbb{F}_{q})\rightarrow W_{n+j}\end{cases}$ (4)
be the vertical coordinate projections, respectively, defined by
$\begin{cases}\pi_{j}(x,t)=(\check{x}_{j},t+\frac{1}{2}x_{j}x_{n+j}),\\ \pi_{n+j}(x,t)=(\check{x}_{n+j},t-\frac{1}{2}x_{j}x_{n+j}).\end{cases}$ (5)
A direct computation shows that for each $j\in\{1,\ldots,2n\}$, an element $(x,t)$ of $\mathbb{H}^{n}(\mathbb{F}_{q})$ can be written
$(x,t)=\pi_{j}(x,t)\cdot(x_{j}e_{j},0),$ (6)
where the dot denotes the group multiplication on $\mathbb{H}^{n}(\mathbb{F}_{q})$.
The fiber of $\pi_{j}$ at a point $(u,\tau)$ in $W_{j}$ is a line in the affine space $\mathbb{F}_{q}^{2n+1}$ passing through $(u,\tau)$. More precisely, for $j\in\{1,\ldots,n\}$, the fiber of $\pi_{j}$ at $(u,\tau)\in W_{j}$ (so $u_{j}=0$) is given by
$\pi_{j}^{-1}\{(u,\tau)\}$ $=\big\{\,\big(u+se_{j},\;\tau-\tfrac{1}{2}\,u_{n+j}\,s\big):\;s\in\mathbb{F}_{q}\,\big\}=(u,\tau)\cdot L_{j},$
and the fiber of $\pi_{n+j}$ at $(u,\tau)\in W_{n+j}$ (so $u_{n+j}=0$) is given by
$\pi_{n+j}^{-1}\{(u,\tau)\}$ $=\big\{\,\big(u+se_{n+j},\;\tau+\tfrac{1}{2}\,u_{j}\,s\big):\;s\in\mathbb{F}_{q}\,\big\}=(u,\tau)\cdot L_{n+j}.$
That is, for each $j\in\{1,\ldots,2n\}$, the fibers of $\pi_{j}$ are exactly cosets of $W_{j}^{\perp}=L_{j}$.
Given a set $K$ in $\mathbb{H}^{n}(\mathbb{F}_{q})$, in this paper, we are interested in the Loomis–Whitney type inequality associated with $K$, namely, bounding the size of $K$ in terms of the sizes of its projections, and its generalizations.
For any set $S$ (in most cases, $S=\mathbb{F}_{q}^{2n})$, denote by $\mathscr{F}(S)$ the set of all $\mathbb{R}$-valued functions on $S$. Note that $\mathscr{F}(S)$ is a vector space over $\mathbb{R}.$ For each $1\leq r\leq\infty$, the $L^{r}$-norm on $\mathscr{F}(\mathbb{F}_{q}^{2n})$ is defined as
$\|f\|_{L^{r}(\mathbb{F}_{q}^{2n},dx)}=\left(\frac{1}{q^{2n}}\sum_{x\in\mathbb{F}_{q}^{2n}}|f(x)|^{r}\right)^{\frac{1}{r}}$
for $1\leq r<\infty$, and for $r=\infty$
$\|f\|_{L^{\infty}(\mathbb{F}_{q}^{2n},dx)}=\max\limits_{x\in\mathbb{F}_{q}^{2n}}|f(x)|.$
Throughout, for each $j\in\{1,\ldots,2n\}$, we shall identify $W_{j}$ with $\mathbb{F}_{q}^{2n}$ via the map $\phi_{j}:W_{j}\stackrel{{\scriptstyle\sim}}{{\rightarrow}}\mathbb{F}_{q}^{2n}$ defined by $(\check{x}_{j},t)\mapsto(\widehat{x}_{j},t)$. For simplicity, for a function $f$ on $\mathbb{F}_{q}^{2n}$ and $(x,t)\in\mathbb{H}^{n}(\mathbb{F}_{q})$, we write $f(\pi_{j}(x,t))$ for $f(\phi_{j}(\pi_{j}(x,t)))$.
Definition 1.1.
Let $n$ be a positive integer. For extended real numbers $1\leq u_{1},u_{2},\ldots,u_{2n}\leq\infty$, define $LW(u_{1},u_{2},\ldots,u_{2n})$ as the smallest constant not depending on $q$, if it exists, such that the inequality
$\frac{1}{q^{2n+1}}\sum_{(x,t)\in\mathbb{H}^{n}(\mathbb{F}_{q})}\prod_{j=1}^{2n}f_{j}(\pi_{j}(x,t))\leq LW(u_{1},u_{2},\ldots,u_{2n})\prod_{j=1}^{2n}\|f_{j}\|_{L^{u_{j}}(\mathbb{F}_{q}^{2n},dx)}$ (7)
holds for all nonnegative $f_{1},f_{2},\ldots,f_{2n}\in\mathscr{F}(\mathbb{F}_{q}^{2n})$. If it does not exist, then we set $LW(u_{1},u_{2},\ldots,u_{2n})=\infty.$
Although our main results are stated as multilinear functional inequalities, they immediately imply corresponding set/projection statements. When $f_{1}=\cdots=f_{2n}$ are all the indicator functions of a subset $K$ of $\mathbb{H}^{n}(\mathbb{F}_{q})$, then the inequality (7) gives a Loomis–Whitney type inequality.
We first present a result in the case $n=1$.
Theorem 1.2.
For $1\leq u_{1},u_{2}\leq\infty,$ the following holds.
$LW(u_{1},u_{2})\lesssim 1\quad\mbox{if and only if}\quad\frac{1}{u_{1}}+\frac{2}{u_{2}}\leq 2\quad\mathrm{and}\quad\frac{2}{u_{1}}+\frac{1}{u_{2}}\leq 2.$
In this paper, we write $X\lesssim Y$ to mean that there exists some constant $C>0$ such that $X\leq CY$.
Corollary 1.3.
For any nonnegative $f_{1},f_{2}\in\mathscr{F}(\mathbb{F}_{q}^{2})$,
$\frac{1}{q^{3}}\sum_{(x,t)\in\mathbb{H}^{1}(\mathbb{F}_{q})}f_{1}(\pi_{1}(x,t))\,f_{2}(\pi_{2}(x,t))\;\lesssim\;\|f_{1}\|_{L^{\frac{3}{2}}(\mathbb{F}_{q}^{2},dx)}\,\|f_{2}\|_{L^{\frac{3}{2}}(\mathbb{F}_{q}^{2},dx)}.$
The endpoint exponent $\frac{3}{2}$ is optimal for general functions $f_{1},f_{2}$ by Theorem 1.2. On the level of sets, Corollary 1.3 implies the bound
$|K|\lesssim q^{\frac{1}{3}}\,|\pi_{1}(K)|^{\frac{2}{3}}\,|\pi_{2}(K)|^{\frac{2}{3}},\qquad K\subset\mathbb{H}^{1}(\mathbb{F}_{q}),$ (8)
and this is sharp up to absolute constants. Indeed, fixing $t_{0}\in\mathbb{F}_{q}$ and taking
$K:=\{(x_{1},0,t_{0})\in\mathbb{H}^{1}(\mathbb{F}_{q})\colon x_{1}\in\mathbb{F}_{q}\},$
we have $|K|=q$, $|\pi_{1}(K)|=1$, and $|\pi_{2}(K)|=q$, so (8) is attained.
This construction can be extended to higher dimensions as follows. Let
$K:=\{(x,t_{0})\in\mathbb{H}^{n}(\mathbb{F}_{q}):x_{n+1}=\cdots=x_{2n}=0\}.$
Then, $|K|=q^{n}$, and
$|\pi_{j}(K)|=\begin{cases}q^{n-1},&1\leq j\leq n,\\[2.0pt] q^{n},&n+1\leq j\leq 2n.\end{cases}$
This example naturally suggests the critical symmetric exponent in higher dimensions. Suppose one seeks a uniform estimate of the form
$\frac{1}{q^{2n+1}}\sum_{(x,t)\in\mathbb{H}^{n}(\mathbb{F}_{q})}\prod_{j=1}^{2n}f_{j}(\pi_{j}(x,t))\;\lesssim\;\prod_{j=1}^{2n}\|f_{j}\|_{L^{u}(\mathbb{F}_{q}^{2n},dx)}$ (9)
with a single exponent $u$ for all inputs. Applying (9) to indicator functions $f_{j}=\mathbf{1}_{\pi_{j}(K)}$ gives
$\frac{|K|}{q^{2n+1}}\;\lesssim\;\prod_{j=1}^{2n}\Bigg(\frac{|\pi_{j}(K)|}{q^{2n}}\Bigg)^{1/u}.$
Inserting the above values of $|K|$ and $|\pi_{j}(K)|$ forces
$q^{n}\;\lesssim\;q^{\,2n+1-\frac{2n^{2}+n}{u}},\qquad\text{hence}\qquad u\;\geq\;\frac{n(2n+1)}{n+1}.$
Thus, $\frac{n(2n+1)}{n+1}$ is the smallest possible symmetric exponent one can expect when $n>1$.
When $n>1$, the full description of the admissible region of $(u_{1},\dots,u_{2n})$ for which $LW(u_{1},\dots,u_{2n})\lesssim 1$ holds is substantially more intricate than in the planar case $n=1$. Rather than attempting a complete characterization, in this paper, we focus on the symmetric critical value $u=\frac{n(2n+1)}{n+1}$, for which the inequality is established in the next theorem.
Theorem 1.4.
For any nonnegative $f_{1},\ldots,f_{2n}\in\mathscr{F}(\mathbb{F}_{q}^{2n})$, we have
$\frac{1}{q^{2n+1}}\sum_{(x,t)\in\mathbb{H}^{n}(\mathbb{F}_{q})}\prod_{j=1}^{2n}f_{j}(\pi_{j}(x,t))\lesssim\prod_{j=1}^{2n}\|f_{j}\|_{L^{\frac{n(2n+1)}{n+1}}(\mathbb{F}_{q}^{2n},dx)}.$ (10)
The proof of the corresponding 2n2n–multilinear inequality is not a formal iteration of the case $n=1$. It is inspired by the induction (on dimension) scheme developed in the continuous setting [7], but requires nontrivial adaptations to the finite field setting and to obtaining constants uniform in $q$. It has two main components. First, we establish a family of asymmetric estimates (Theorem 4.1) in which one symplectic pair of inputs is measured in $L^{\frac{2n+1}{2}}$ while the remaining inputs are measured in $L^{2n+1}$. These bounds are proved by a delicate induction on $n$ that repeatedly freezes one symplectic coordinate pair and rewrites the nn–dimensional form as an averaged (n1)(n-1)–dimensional Heisenberg Loomis–Whitney form. Second, we combine the resulting family of asymmetric bounds through a multilinear interpolation scheme to reach the symmetric estimate at the critical exponent $u=\frac{n(2n+1)}{n+1}$.
As an application of Theorem 1.4, we obtain the following.
Corollary 1.5.
For $K\subset\mathbb{H}^{n}(\mathbb{F}_{q})$, we have
$|K|\lesssim q^{\frac{1}{2n+1}}\prod_{j=1}^{2n}|\pi_{j}(K)|^{\frac{n+1}{n(2n+1)}}.$ (11)
In particular, for $K\subset\mathbb{H}^{1}(\mathbb{F}_{q})$,
$\max\{|\pi_{1}(K)|,|\pi_{2}(K)|\}\gtrsim|K|^{\frac{3}{4}}q^{-\frac{1}{4}}.$
It follows from the example above that this corollary is sharp and we cannot replace the factor $q^{\frac{1}{2n+1}}$ by $|\pi_{v}(K)|^{\frac{1}{2n+1}}$, where $\pi_{v}(K)$ is the image of $K$ under the (vertical) projection $\pi_{v}:\mathbb{H}^{n}(\mathbb{F}_{q})\rightarrow\mathbb{F}_{q}$ defined by $(x,t)\mapsto t$.
When $n=1$ and $|K|>q$, by using the point–line incidence theorem due to Vinh [20], we are able to provide a better lower bound for the quantity $\max\{|\pi_{1}(K)|,|\pi_{2}(K)|\}$.
Theorem 1.6.
For $K\subset\mathbb{H}^{1}(\mathbb{F}_{q})$, we have
$|K|\lesssim\frac{|\pi_{1}(K)|\,|\pi_{2}(K)|}{q}+q^{\frac{1}{2}}\sqrt{|\pi_{1}(K)|\,|\pi_{2}(K)|}.$
In particular,
$\max\{|\pi_{1}(K)|,|\pi_{2}(K)|\}\gtrsim\min\left\{|K|^{\frac{1}{2}}q^{\frac{1}{2}},~|K|q^{-\frac{1}{2}}\right\}.$
This theorem is also sharp in its range. To see this, let $A,B\subset\mathbb{F}_{q}$ with $|A|=|B|=m$, and put $T=\mathbb{F}_{q}$. Define
$K:=A\times B\times T\;\subset\;\mathbb{H}^{1}(\mathbb{F}_{q}).$
Then
$\pi_{1}(K)=B\times\mathbb{F}_{q},\quad|\pi_{1}(K)|=mq,\qquad\pi_{2}(K)=A\times\mathbb{F}_{q},\quad|\pi_{2}(K)|=mq,$
and
$|K|=|A|\,|B|\,|T|=m^{2}q.$
Consequently,
$\frac{|\pi_{1}(K)|\,|\pi_{2}(K)|}{q}\;+\;q^{\frac{1}{2}}\sqrt{|\pi_{1}(K)|\,|\pi_{2}(K)|}\;=\;m^{2}q\;+\;mq^{\frac{3}{2}},$
which matches $|K|$ up to absolute constants.
2 Subgroups of the Heisenberg group and projections
This section is devoted to the study of subgroups $G$ of the Heisenberg group $\mathbb{H}^{n}(\mathbb{F}_{q})$ and the analysis of the projections onto relevant subgroups. The structural statements developed here will be used throughout the paper and provide tools for subsequent work.
2.1 Basics on subgroups of $\mathbb{H}^{n}(\mathbb{F}_{q})$
When viewing a vector space $V$ over $\mathbb{F}_{q}$ as a group, we consider only its additive group structure, forgetting scalar multiplication. Trivially, a subspace of the vector space $V$ is a subgroup of $V$. If the field $\mathbb{F}_{q}$ is a prime field, i.e., $q=p$ for some prime $p$, then the converse is also true. Indeed, if $W$ is a subgroup of $V,$ then we can give a scalar multiplication on $W$ that makes $W$ into a subspace of $V.$ To see this, we write $\mathbb{F}_{p}=\{0,1,\ldots,p-1\}.$ For $t\in\mathbb{F}_{p}$ and $w\in W$, define $tw:=w+\cdots+w$ ($t$-times). We easily see that this coincides with the scalar multiplication on $W$ induced from $V$, and so $W$ is a subspace of $V.$ If $\mathbb{F}_{q}$ is not a prime field, then there is a subgroup $W$ of $V$ that is not a subspace. For example, if $q=p^{r}$ for some $r\geq 2,$ the subfield $\mathbb{F}_{p}$ is a subgroup of the (one dimensional) vector space $\mathbb{F}_{q},$ but not a subspace of $\mathbb{F}_{q}$ (over $\mathbb{F}_{q}$). Over a general $\mathbb{F}_{q}$ with $q=p^{r},$ in order to obtain all subgroups of a vector space $V$, it is enough to view $V$ as a vector space over $\mathbb{F}_{p}$ and find all subspaces of $V$ over $\mathbb{F}_{p}.$ Similarly, linear maps between two vector spaces over a prime field are precisely group homomorphisms between the corresponding underlying groups. For a subspace $W$ of the vector space $\mathbb{F}_{q}^{m},$ let $W^{\perp}$ be the subspace of $\mathbb{F}_{q}^{m}$ defined as $W^{\perp}=\{v\in\mathbb{F}_{q}^{m}:\langle v,w\rangle=0\,\,\mathrm{for\,\,\,all}\,w\in W\}$, where $\langle\,$$\rangle$ is the standard inner product on $\mathbb{F}_{q}^{m}.$ We may call $W^{\perp}$ the orthogonal complement of $W$ in $\mathbb{F}_{q}^{m}$ with respect to $\langle\,$$\rangle$.
The quadratic term in the group multiplication (2) on $\mathbb{H}^{n}(\mathbb{F}_{q})$ plays an important role in understanding $\mathbb{H}^{n}(\mathbb{F}_{q})$ and its subgroups. So we treat it especially. For $x,x^{\prime}\in\mathbb{F}_{q}^{2n}$, define
$\omega(x,x^{\prime}):=\sum_{i=1}^{n}(x_{i}x_{n+i}^{\prime}-x_{i}^{\prime}x_{n+i}).$
Then $\omega$ is a symplectic form, i.e., a nondegenerate skew-symmetric bilinear form on $\mathbb{F}_{q}^{2n}$. Note that $\omega$ is determined by the skew symmetric matrix $A_{n}$, i.e., $\omega(x,x^{\prime})=x\cdot A_{n}\cdot(x^{\prime})^{t},$ where
$A_{n}:=\begin{bmatrix}0&I_{n}\\ -I_{n}&0\\ \end{bmatrix}$,
and $(\,)^{t}$ stands for the transpose of a matrix.
Definition 2.1.
A subspace $S$ of $\mathbb{F}_{q}^{2n}$ is called isotropic if $\omega(x,x^{\prime})=0$ for any two vectors $x,x^{\prime}\in S$.
Note that by linear algebra, maximal isotropic subspaces of $\mathbb{F}_{q}^{2n}$ have dimension $n$, and all subspaces of dimension one are isotropic.
Let $\pi_{h}:\mathbb{H}^{n}(\mathbb{F}_{q})\rightarrow\mathbb{F}_{q}^{2n}$ be the projection defined by $\pi_{h}(x,t)=x$. Then $\pi_{h}$ is a group homomorphism and so if $G$ is a subgroup of $\mathbb{H}^{n}(\mathbb{F}_{q}),$ then the image $\pi_{h}(G)$ is a subgroup of $\mathbb{F}_{q}^{2n}.$ Recall that $\pi_{v}:\mathbb{H}^{n}(\mathbb{F}_{q})\rightarrow\mathbb{F}_{q}$ is the projection defined by $\pi_{v}(x,t)=t$. Note that $\pi_{v}$ is not a group homomorphism, but its restriction to the ‘vertical line’ $\{(0,t)\in\mathbb{H}^{n}(\mathbb{F}_{q})\,|\,t\in\mathbb{F}_{q}\}$ is a group isomorphism.
Define an action of the group $\mathbb{F}_{q}^{*}$ on $\mathbb{H}^{n}(\mathbb{F}_{q})$ by $s\cdot(x,t)=(sx,s^{2}t)$ for $s\in\mathbb{F}_{q}^{*}$ and $(x,t)\in\mathbb{H}^{n}(\mathbb{F}_{q}).$ This action shall be called the dilation action, following [2].
Definition 2.2.
A subgroup $G$ of $\mathbb{H}^{n}(\mathbb{F}_{q})$ is called homogeneous if $G$ is invariant under the dilation action of $\mathbb{F}_{q}^{*}$ on $\mathbb{H}^{n}(\mathbb{F}_{q})$.
Recall that the orbit of an element $(x,t)\in\mathbb{H}^{n}(\mathbb{F}_{q})$ under the dilation action is given by
$\mathcal{O}_{\mathbb{F}_{q}^{*}}(x,t):=\{s\cdot(x,t)\in\mathbb{H}^{n}(\mathbb{F}_{q})\,|\,s\in\mathbb{F}_{q}^{*}\}.$
If $G$ is any homogeneous subgroup of $\mathbb{H}^{n}(\mathbb{F}_{q})$ containing $(x,t)$, then $G$ necessarily includes $\mathcal{O}_{\mathbb{F}_{q}^{*}}(x,t)$.
Lemma 2.3.
For $(x,t)\in\mathbb{H}^{n}(\mathbb{F}_{q})$, the cardinality of $\mathcal{O}_{\mathbb{F}_{q}^{*}}(x,t)$ is given as follows.
  • (i)
    If $x\neq 0$, then $|\mathcal{O}_{\mathbb{F}_{q}^{*}}(x,t)|=q-1.$
  • (ii)
    If $x=0$ and $t\neq 0$, then $|\mathcal{O}_{\mathbb{F}_{q}^{*}}(x,t)|=\frac{q-1}{2}.$
  • (iii)
    If $x=0$ and $t=0$, then $|\mathcal{O}_{\mathbb{F}_{q}^{*}}(x,t)|=1.$
Proof.
It can be easily verified that the isotropy groups of $(x,t)$ for the cases (i), (ii), (iii)
$\mathrm{ISO}_{\mathbb{F}_{q}^{*}}(x,t)=\{s\in\mathbb{F}_{q}^{*}\,|\,s\cdot(x,t)=(x,t)\}$
are $\{1\},$ $\mathbb{Z}_{2}=\{-1,1\}$ and $\mathbb{F}_{q}^{*}$ , respectively. Then, the lemma is immediate from the formula
$|\mathcal{O}_{\mathbb{F}_{q}^{*}}(x,t)|\cdot|\mathrm{ISO}_{\mathbb{F}_{q}^{*}}(x,t)|=|\mathbb{F}_{q}^{*}|=q-1.$
Now we will identify and count all subgroups $G$ of $\mathbb{H}^{n}$ for two special cases, namely the case when $\mathbb{F}_{q}$ is a prime field, and the case where subgroups $G$ are homogeneous.
First assume that $\mathbb{F}_{q}$ is a prime field, say, $q=p$ for some prime $p.$ In this case we write $\mathbb{F}_{p}$ for $\mathbb{F}_{q}$. First fix a subgroup (equivalently, a subspace) $S$ of $\mathbb{F}_{p}^{2n}.$ If $G$ is a subgroup of $\mathbb{H}^{n}(\mathbb{F}_{p})$ such that $\pi_{h}(G)=S$, by cardinality reason for finite groups, there are only two possibilities
$|G|=|S|,$
$|G|=|S|p.$
If $|G|=|S|p,$ then necessarily $G=S\times\mathbb{F}_{p}.$ If $|G|=|S|,$ then $G$ is a graph $\Gamma_{S}(\rho)$ of a map $\rho:S\rightarrow\mathbb{F}_{p}$,
$\Gamma_{S}(\rho):=\{(x,t)\subset\mathbb{H}^{n}(\mathbb{F}_{p})\,|\,x\in S,t=\rho(x)\}.$
But since $G=\Gamma_{S}(\rho)$ is a group, i.e., closed under the induced multiplication, $S$ should be isotropic and $\rho$ is a group homomorphism. Summarizing, we have
Proposition 2.4.
If $G$ is a subgroup of $\mathbb{H}^{n}(\mathbb{F}_{p})$ for a prime $p$, then $G$ is one of following two types
  • (a)
    $G=S\times\mathbb{F}_{p}$ for a subspace $S$ of $\mathbb{F}_{p}^{2n}$.
  • (b)
    $G=\Gamma_{S}(\rho)$ for a group homomorphism (equivalently, a linear map) $\rho:S\rightarrow\mathbb{F}_{p}$, where $S$ is an isotropic subspace of $\mathbb{F}_{p}^{2n}$.
Next, let us treat homogeneous subgroups of $\mathbb{H}^{n}(\mathbb{F}_{q})$. If $G$ is a homogeneous subgroup of $\mathbb{H}^{n}(\mathbb{F}_{q})$, then $\pi_{h}(G)$ is a subspace of $\mathbb{F}_{q}^{2n}$ (over $\mathbb{F}_{q}$). So, we fix a subspace $S$ of $\mathbb{F}_{q}^{2n}$, then find all subgroups $G$ of $\mathbb{H}^{n}(\mathbb{F}_{q})$ such that $\pi_{h}(G)=S.$ Assume that $q=p^{r}$ for some prime $p$ and a positive integer $r$. The projection $\pi_{h}:\mathbb{H}^{n}(\mathbb{F}_{q})\rightarrow\mathbb{F}_{q}^{2n}$ induces a group homomorphism $\pi_{h}|_{G}:$ $G\rightarrow S.$ Then, elements of the kernel $K$ of $\pi_{h}|_{G}$ are of the form $(0,t)$ for $t\in\mathbb{F}_{q}$. Also, note that the restriction $\pi_{v}|_{K}:K\rightarrow\mathbb{F}_{q}$ embeds $K$ into $\mathbb{F}_{q}$ as a subgroup. Thus, $|K|$ divides $q=p^{r}$, and so $|K|=p^{\ell}$ for some $\ell\leq r.$ Firstly, assume that $K$ is nontrivial, i.e., contains an element $(0,t)$ with $t\neq 0.$ Since $K$ is homogeneous, $K$ includes $\mathcal{O}_{\mathbb{F}_{q}^{*}}(0,t)$. So, by Lemma 2.3 (ii), we obtain $|K|\geq\frac{q-1}{2}=\frac{p^{r}-1}{2}$. This forces $|K|=p^{r}=q$ since $p$ is odd. (Note that if $\ell\leq r-1$, then it would be that $|K|<\frac{q-1}{2}.$) Since $|G|=|S||K|=|S|q,$ in this case we have $G=S\times\mathbb{F}_{q}.$ Secondly, assume that $K=\{(0,0)\}$ is trivial. Then $\pi_{h}|_{G}:G\rightarrow S$ is an isomorphism, and so, as in Proposition 2.4, $S$ is isotropic and $G=\Gamma_{S}(\rho)$ for some group homomorphism(equivalently, a linear map over $\mathbb{F}_{p}$) $\rho:S\rightarrow\mathbb{F}_{q}.$ We claim that $\rho$ is trivial. To show this, let $(x,\rho(x))\in\Gamma_{S}(\rho)=G$ with $0\neq x\in\mathbb{F}_{q}^{2n}$. Then by the homogeneity condition, for any $s\in\mathbb{F}_{q}^{*}$, we have
$s\cdot(x,\rho(x))=(sx,s^{2}\rho(x))=(sx,\rho(sx)).$ (12)
But, in particular, if $s\in\mathbb{F}_{p}^{*}$, then $\rho(sx)=s\rho(x)$ since $\rho$ is linear over $\mathbb{F}_{p}.$ By this fact and (12), we obtain the equality $s^{2}\rho(x)=s\rho(x)$ for any $s\in\mathbb{F}_{p}^{*}$. This implies that $\rho(x)=0$ for all nonzero $x\in\mathbb{F}_{q}^{2n}$, i.e., $\rho$ is trivial. Therefore, $G=\Gamma_{S}(\rho)=S\times\{0\}.$ In conclusion, we have proved
Proposition 2.5.
If $G$ is a homogeneous subgroup of $\mathbb{H}^{n}(\mathbb{F}_{q})$, then $G$ is one of the following two cases
  • (a)
    $G=S\times\mathbb{F}_{q}$ for a subspace $S$ of $\mathbb{F}_{q}^{2n}$.
  • (b)
    $G=S\times\{0\}$ for an isotropic subspace $S$ of $\mathbb{F}_{q}^{2n}$.
2.2 Orthogonal complement of a homogeneous subgroup of $\mathbb{H}^{n}(\mathbb{F}_{q})$
By Proposition 2.5, homogeneous subgroups $G$ of $\mathbb{H}^{n}(\mathbb{F}_{q})$ are some subspaces of $\mathbb{F}_{q}^{2n+1}$ via the identification $\mathbb{H}^{n}(\mathbb{F}_{q})=\mathbb{F}_{q}^{2n+1}$. By abuse of notation, through this identification, we take $G^{\perp}$ for homogeneous subgroups $G$ of $\mathbb{H}^{n}(\mathbb{F}_{q})$. If $G=S\times\{0\}$ for an isotropic $S$ is a homogeneous subgroup, then we have $G^{\perp}=S^{\perp}\times\mathbb{F}_{q}$, which is necessarily a (vertical) homogeneous subgroup of $\mathbb{H}^{n}(\mathbb{F}_{q})$. If $G=S\times\mathbb{F}_{q}$, where $S$ is an arbitrary subspace of $\mathbb{F}_{q}^{2n}$, then $G^{\perp}=S^{\perp}\times\{0\}$. In this case $G^{\perp}$ is not necessarily a homogeneous subgroup of $\mathbb{H}^{n}(\mathbb{F}_{q})$. For example, if $\mathrm{dim}\,S\leq n-1$, then $\mathrm{dim}\,S^{\perp}\geq n+1$ and so $S$ is not isotropic by the dimension reason. Thus $G^{\perp}=S^{\perp}\times\{0\}$ is not a homogeneous subgroup of $\mathbb{H}^{n}(\mathbb{F}_{q})$. (In fact, it is not even a subgroup of $\mathbb{H}^{n}(\mathbb{F}_{q})$.)
2.3 Counting subgroups of $\mathbb{H}^{n}(\mathbb{F}_{q})$
Now we count subgroups of $\mathbb{H}^{n}(\mathbb{F}_{q})$ for the two special cases in the above. For $k=0,\ldots,2n$, let $\mathrm{Gr}(k,2n)$ be the variety of $k$ dimensional subspaces in $\mathbb{F}_{q}^{2n}.$ For $k=0,\ldots,n$, let $\mathrm{IG}(k,2n)$ be the variety of $k$ dimensional isotropic subspaces in $\mathbb{F}_{q}^{2n}$.
First we count subgroups of $\mathbb{H}^{n}(\mathbb{F}_{p})$ for a prime $p$, using Proposition 2.4. For a subspace $S$ of $\mathbb{F}_{p}^{2n}$ of dimension $k$ ($|S|=p^{k}$), let n(S)n(S)(resp. $n(S)^{\prime}$) be the number of subgroups $G$ of $\mathbb{H}^{n}(\mathbb{F}_{p})$ such that $\pi_{h}(G)=S$ and $|G|=p^{k}$ (resp. $|G|=p^{k+1}$). Then, by Proposition 2.4, the number of all subgroups of $\mathbb{H}^{n}(\mathbb{F}_{p})$ equals
$\sum_{k=0}^{2n}\sum_{S\in\mathrm{Gr}(k,2n)}|n(S)^{\prime}|+\sum_{k=0}^{n}\sum_{S\in\mathrm{IG}(k,2n)}|n(S)|.$
We easily see that $n(S)^{\prime}=1$ for each $S\in\mathrm{Gr}(k,2n)$; and $n(S)=p^{k}$ for each $S\in\mathrm{IG}(k,2n)$ because there are $p^{k}$ group homomorphisms (linear maps) from $S$ to $\mathbb{F}_{p}$, identified with $1\times k$ matrices with entries in $\mathbb{F}_{p}$.
This number equals
$\sum_{k=0}^{2n}|\mathrm{Gr}(k,2n)|+\sum_{k=0}^{n}p^{k}|\mathrm{IG}(k,2n)|.$ (13)
Next, we count homogeneous subgroups of $\mathbb{H}^{n}(\mathbb{F}_{q})$, using Proposition 2.5. Fix a subspace $S$ of $\mathbb{F}_{q}^{2n}$ of dimension $k$ (so $|S|=q^{k}$). For such $S$, let $m(S)$ (resp. $m(S)^{\prime}$) be the number of homogeneous subgroups $G$ with $\pi_{h}(G)=S$ and $|G|=q^{k}$ (resp. $|G|=q^{k+1}$). Then, by Proposition 2.5, the number of all homogeneous subgroups of $\mathbb{H}^{n}(\mathbb{F}_{q})$ equals
$\sum_{k=0}^{2n}\sum_{S\in\mathrm{Gr}(k,2n)}|m(S)^{\prime}|+\sum_{k=0}^{n}\sum_{S\in\mathrm{IG}(k,2n)}|m(S)|.$
Since $m(S)^{\prime}=1$ for each $S\in\mathrm{Gr}(k,2n)$; and $m(S)=1$ for each $S\in\mathrm{IG}(k,2n)$, the number of all homogeneous subgroups of $\mathbb{H}^{n}(\mathbb{F}_{q})$ equals
$\sum_{k=0}^{2n}|\mathrm{Gr}(k,2n)|+\sum_{k=0}^{n}|\mathrm{IG}(k,2n)|.$ (14)
Now, by (13) and (14), enumerating subgroups of $\mathbb{H}^{n}(\mathbb{F}_{q})$ is reduced to counting elements of $|\mathrm{Gr}(k,2n)|$ and $|\mathrm{IG}(k,2n)|$. Let us give counting formulas for elements of these varieties. For an indeterminate $q$ and an integer $m>0$, let $m]_{q}$ be a polynomial in $q$
$m]_{q}:=\sum_{i=0}^{m-1}q^{i}=\frac{1-q^{m}}{1-q},$
and set
$m]_{q}!:=[m]_{q}[m-1]_{q}\cdots[2]_{q}[1]_{q}.$
Then, the following formulas can be found in the literature, for example, in [5, 19].
Proposition 2.6.
The cardinalities of $\mathrm{Gr}(k,2n)$ and $\mathrm{IG}(k,2n)$, respectively, are given by
$|\mathrm{Gr}(k,2n)|=\frac{[2n]_{q}!}{[k]_{q}![(2n-k)]_{q}!},$
$|\mathrm{IG}(k,2n)|=\frac{[n]_{q}!}{[k]_{q}![n-k]_{q}!}\cdot\prod_{i=n-k+1}^{n}\left(q^{i}+1\right).$
In Proposition 2.5, we dealt with only homogeneous subgroups of $\mathbb{H}^{n}(\mathbb{F}_{q})$. Using the idea in Propositions 2.4 and 2.5 we can still construct some subgroups of $\mathbb{H}^{n}(\mathbb{F}_{q})$. For example, fix a subgroup $S$ of $\mathbb{F}_{q}^{2n}$ so that $S$ is a subspace over $\mathbb{F}_{p}$. Let $F(S)$ be the smallest subfield of $\mathbb{F}_{q}$ containing all components $x_{i}\in\mathbb{F}_{q}$ of $x=(x_{1},\ldots,x_{2n})$, where $x$ runs through $S$. Then for any subfield $F$ of $\mathbb{F}_{q}$ with $F(S)\subset F$, the product $S\times F$ is a subgroup of $\mathbb{H}^{n}(\mathbb{F}_{q})$. Also, if $S$ is isotropic, then as in Proposition 2.4, we can obtain a subgroup $G$ of $\mathbb{H}^{n}(\mathbb{F}_{q})$ as a graph of a group homomorphism. The group thus constructed is of the non-product type. In case $S$ is not isotropic, we could not find a subgroup $G$ of the non-product type, with $\pi_{h}(G)=S$, and we do not know whether there exists such a subgroup $G$. It would be interesting to investigate existence of such subgroups of the non-product type, with $\pi_{h}(G)=S$ for a non-isotropic subgroup $S$, or, more generally, to classify and count all subgroups of $\mathbb{H}^{n}(\mathbb{F}_{q})$.
In conclusion, the coordinate subgroups $W_{1},\dots,W_{2n}$ discussed here form a homogeneous family that will be central to our main results. A detailed analysis of other homogeneous subgroups will be reserved for future work, to maintain focus and clarity in the current discussion.
3 Proof of Theorem 1.2
Define a bilinear operator (form) $\mathcal{L}$ on $\mathscr{F}(\mathbb{F}_{q}^{2})$ by
$\mathcal{L}(f_{1},f_{2}):=\frac{1}{q^{3}}\sum_{x=(x_{1},x_{2}),y=(y_{1},y_{2})\in\mathbb{F}_{q}^{2}}1_{x_{1}\cdot y_{1}+y_{2}=x_{2}}f_{1}(x)f_{2}(y),$ (15)
for any $f_{1},f_{2}\in\mathscr{F}(\mathbb{F}_{q}^{2}).$ Let $\langle\,$$\rangle$ be the ‘standard’ inner product on $\mathscr{F}(\mathbb{F}_{q}^{2})$, i.e.,
$\langle f_{1},f_{2}\rangle=\frac{1}{q^{2}}\sum_{x=(x_{1},x_{2})\in\mathbb{F}_{q}^{2}}f_{1}(x)f_{2}(x).$
Then we can write $\mathcal{L}(f_{1},f_{2})=\langle f_{1},Af_{2}\rangle$, where $A$ is an endomorphism on $\mathscr{F}(\mathbb{F}_{q}^{2})$ defined by
$\left(Af_{2}\right)(x):=\frac{1}{q}\sum_{y=(y_{1},y_{2})\in\mathbb{F}_{q}^{2}}1_{x_{1}y_{1}+y_{2}=x_{2}}f_{2}(y)$
for $f\in\mathscr{F}(\mathbb{F}_{q}^{2})$. Note that the bilinear operator $\mathcal{L}$ is not symmetric.
For $1\leq r,s\leq\infty,$ denote by $A(s\to r)$ the smallest number, if it exists, such that the inequality
$\|Af\|_{L^{r}(\mathbb{F}^{2}_{q},dx)}\leq A(s\to r)\|f\|_{L^{s}(\mathbb{F}^{2}_{q},dx)}$
holds for all nonnegative $f\in\mathscr{F}(\mathbb{F}_{q}^{2}).$
To prove Theorem 1.2, we introduce a constant $L(u_{1},u_{2})$ as an alternative to the constant $LW(u_{1},u_{2}).$
Definition 3.1.
For $1\leq u_{1},u_{2}\leq\infty,$ define $L(u_{1},u_{2})$ as the smallest positive constant, if it exists, such that the following estimate holds for all nonnegative $f_{1},f_{2}\in\mathscr{F}(\mathbb{F}_{q}^{2})$
$\mathcal{L}(f_{1},f_{2})\leq L(u_{1},u_{2})\|f_{1}\|_{L^{u_{1}}(\mathbb{F}^{2}_{q},dx)}\cdot\|f_{2}\|_{L^{u_{2}}(\mathbb{F}^{2}_{q},dx)},$ (16)
and if it does not exist, we set $L(u_{1},u_{2})=\infty.$
Recall that for $1\leq u\leq\infty$, the Hölder’s conjugate of $u$ is defined by $\frac{1}{u}+\frac{1}{u^{\prime}}=1$.
Lemma 3.2.
If $1\leq u_{1},u_{2}\leq\infty$ satisfy $A(u_{2}\to u_{1}^{\prime})\lesssim 1$, then we have $L(u_{1},u_{2})\lesssim 1.$
Proof.
By Hölder’s inequality and the assumption, we have
$\mathcal{L}(f_{1},f_{2})=\langle f_{1},Af_{2}\rangle\leq\|f_{1}\|_{L^{u_{1}}(\mathbb{F}^{2}_{q},dx)}\,\|Af_{2}\|_{L^{u^{\prime}_{1}}(\mathbb{F}^{2}_{q},dx)}\lesssim\|f_{1}\|_{L^{u_{1}}(\mathbb{F}^{2}_{q},dx)}\,\|f_{2}\|_{L^{u_{2}}(\mathbb{F}^{2}_{q},dx)}.$
Since these inequalities hold for any $f_{1},f_{2}\in\mathscr{F}(\mathbb{F}_{q}^{2})$, the lemma is immediate. ∎
Theorem 3.3.
For $1\leq u_{1},u_{2}\leq\infty,$ the following holds.
$L(u_{1},u_{2})\lesssim 1\quad\mbox{if and only if}\quad\frac{1}{u_{1}}+\frac{2}{u_{2}}\leq 2\quad\mathrm{and}\quad\frac{2}{u_{1}}+\frac{1}{u_{2}}\leq 2.$
Proof.
First, assume $L(u_{1},u_{2})\lesssim 1$. To obtain the first inequality $\tfrac{1}{u_{1}}+\tfrac{2}{u_{2}}\leq 2$, take two test functions
$f_{1}(x_{1},x_{2})=\mathbf{1}_{\{x_{2}=x_{1}+1\}},\quad f_{2}(x_{1},x_{2})=\mathbf{1}_{\{x_{1}=x_{2}=1\}}.$
By simple calculations, we obtain
$\mathcal{L}(f_{1},f_{2})=q^{-2};$
$\|f_{1}\|_{L^{u_{1}}(\mathbb{F}^{2}_{q},dx)}=\Bigl(\frac{q}{q^{2}}\Bigr)^{1/u_{1}}=q^{-1/u_{1}},\qquad\|f_{2}\|_{L^{u_{2}}(\mathbb{F}^{2}_{q},dx)}=q^{-2/u_{2}}.$
The assumption $L(u_{1},u_{2})\lesssim 1$ gives
$q^{-2}\lesssim q^{-1/u_{1}-2/u_{2}},$
which is equivalent to the inequality $\tfrac{1}{u_{1}}+\tfrac{2}{u_{2}}\leq 2$.
The second inequality $\tfrac{2}{u_{1}}+\tfrac{1}{u_{2}}\leq 2$ can be obtained in a similar manner by choosing
$g_{1}(x_{1},x_{2})=\mathbf{1}_{\{x_{1}=x_{2}=1\}},\quad g_{2}(x_{1},x_{2})=\mathbf{1}_{\{x_{2}=1-x_{1}\}}.$
For the other direction, assume that $1\leq u_{1},u_{2}\leq\infty$ satisfy the inequalities
$\frac{1}{u_{1}}+\frac{2}{u_{2}}\leq 2\quad\mathrm{and}\quad\frac{2}{u_{1}}+\frac{1}{u_{2}}\leq 2.$
Let us show that $L(u_{1},u_{2})\lesssim 1$. By the interpolation theorem and the nesting property of the norm, it suffices to establish the estimates on the critical end-points $(\frac{1}{u_{1}},\frac{1}{u_{2}})\in[0,1]\times[0,1]$, which are $(0,1),(1,0),$ and $(\frac{2}{3},\frac{2}{3})$. In other words, it remains to prove the following estimates:
$L(\infty,1)\lesssim 1,~L(1,\infty)\lesssim 1,~L\left(\frac{3}{2},~\frac{3}{2}\right)\lesssim 1.$
But these estimates follow directly from Lemma 3.2 and the following result, due to Koh [13, Theorem 1.1]:
$A(1\to 1)\lesssim 1,~~A(\infty\to\infty)\lesssim 1,~~A\left(\frac{3}{2}\to 3\right)\lesssim 1.$
The following result is a special case of Theorem 3.3, but it is very useful in practice.
Corollary 3.4.
We have $L\left(\frac{3}{2},\frac{3}{2}\right)\lesssim 1.$
Proof.
The proof follows immediately from Theorem 3.3 since
$\frac{1}{u_{1}}+\frac{2}{u_{2}}=\frac{2}{u_{1}}+\frac{1}{u_{2}}=2$
with $u_{1}=u_{2}=\frac{3}{2}$. ∎
Now we can compare two constants $LW(u_{1},u_{2})$ and $L(u_{1},u_{2}).$
Lemma 3.5.
For $1\leq u_{1},u_{2}\leq\infty$, the following holds.
$L(u_{1},u_{2})\lesssim 1\quad\mathrm{if\,\,and\,\,only\,\,if}\quad LW(u_{1},u_{2})\lesssim 1.$
Proof.
To prove the lemma, it is sufficient to show the LHS of the inequality (7) with $n=1$ is equal to LHS, $\mathcal{L}(f_{1},f_{2})$, of the inequality (16). In fact, it follows from (7.1) that
$\frac{1}{q^{3}}\sum_{(x_{1},x_{2},t)\in\mathbb{H}_{q}^{1}}f_{1}(\pi_{1}(x_{1},x_{2},t))f_{2}(\pi_{2}(x_{1},x_{2},t))=\frac{1}{q^{3}}\sum_{x_{1},x_{2},t\in\mathbb{F}_{q}}f_{1}(x_{2},t+x_{1}x_{2})f_{2}(x_{1},t)=\mathcal{L}(f_{1},f_{2}),$
as desired. ∎
In conclusion, Theorem 1.2 follows from Theorem 3.3 and Lemma 3.5
4 Proof of Theorem 1.4
While proving Theorem 1.4, we will heavily use the following result.
Theorem 4.1.
Fix any $k=1,\ldots,n$. Then for any nonnegative $f_{1},\ldots,f_{2n}\in\mathscr{F}(\mathbb{F}_{q}^{2n}),$ the inequality
$\frac{1}{q^{2n+1}}\sum_{(x,t)\in\mathbb{H}^{n}(\mathbb{F}_{q})}\prod_{j=1}^{2n}f_{j}(\pi_{j}(x,t))$
$\lesssim\|f_{k}\|_{L^{\frac{2n+1}{2}}(\mathbb{F}_{q}^{2n},dx)}\|f_{n+k}\|_{L^{\frac{2n+1}{2}}(\mathbb{F}_{q}^{2n},dx)}\prod_{j=1,j\neq k}^{n}\left(\|f_{j}\|_{L^{2n+1}(\mathbb{F}_{q}^{2n},dx)}\|f_{n+j}\|_{L^{2n+1}(\mathbb{F}_{q}^{2n},dx)}\right)$ (17)
holds.
Proof.
We may assume, WLOG, that $k=1$. Denote
$I:=\frac{1}{q^{2n+1}}\sum_{(x,t)\in\mathbb{H}^{n}(\mathbb{F}_{q})}\prod_{j=1}^{2n}f_{j}(\pi_{j}(x,t)).$
Step $1^{\circ}$ (Separate the factor $f_{n}$.) For convenience, for $x=(x_{1},\ldots,x_{2n})\in\mathbb{F}_{q}^{2n}$, write $x=(x^{\prime},y)$, where $x^{\prime}=(x_{1},\ldots,x_{2n-1})\in\mathbb{F}_{q}^{2n-1}$ and $y=x_{2n}\in\mathbb{F}_{q}$, and the change of variables $\tau:=t-\frac{1}{2}x_{n}y$. Then we can write
$I$ $=\frac{1}{q^{2n+1}}\sum_{(x^{\prime},y,t)\in\mathbb{H}_{q}^{n}}f_{2n}(x^{\prime},t-\frac{1}{2}x_{n}y)f_{n}(\widehat{x}_{n},t+\frac{1}{2}x_{n}y)\,\prod_{\begin{subarray}{c}j=1\\ j\neq n,2n\end{subarray}}^{2n}f_{j}\big(\pi_{j}(x^{\prime},y,t)\big)$
$=\frac{1}{q^{2n+1}}\sum_{(x^{\prime},y,\tau)\in\mathbb{F}_{q}^{2n+1}}f_{2n}(x^{\prime},\tau)\,f_{n}(\widehat{x}_{n},\tau+x_{n}y)\prod_{\begin{subarray}{c}j=1\\ j\neq n,2n\end{subarray}}^{2n}f_{j}\big(\pi_{j}(x^{\prime},y,\tau+\frac{1}{2}x_{n}y)\big)$
$=\frac{1}{q^{2n}}\sum_{(x^{\prime},\tau)\in\mathbb{F}_{q}^{2n}}f_{2n}(x^{\prime},\tau)\,\frac{1}{q}\sum_{y\in\mathbb{F}_{q}}f_{n}(\widehat{x}_{n},\tau+x_{n}y)\prod_{\begin{subarray}{c}j=1\\ j\neq n,2n\end{subarray}}^{2n}f_{j}\big(\pi_{j}(x^{\prime},y,\tau+\frac{1}{2}x_{n}y)\big).$ (18)
Note that $f_{2n}$ is independent of $y$. By applying Hölder’s inequality with the conjugate pair of exponents $(2n+1,\frac{2n+1}{2n})$, we obtain
$I\leq$ $\bigg(\frac{1}{q^{2n}}\sum_{(x^{\prime},\tau)\in\mathbb{F}_{q}^{2n}}f_{2n}(x^{\prime},\tau)^{2n+1}\bigg)^{\frac{1}{2n+1}}\cdot J=\|f_{2n}\|_{L^{2n+1}(\mathbb{F}_{q}^{2n},dx)}\cdot J,$
where $J$ captures the contribution of the remaining product
$J:=\left[\frac{1}{q^{2n}}\sum_{(x^{\prime},\tau)\in\mathbb{F}_{q}^{2n}}\bigg(\frac{1}{q}\sum_{y\in\mathbb{F}_{q}}f_{n}(\widehat{x}_{n},\tau+x_{n}y)\prod_{\begin{subarray}{c}j=1\\ j\neq n,2n\end{subarray}}^{2n}f_{j}\big(\pi_{j}(x^{\prime},y,\tau+\frac{1}{2}x_{n}y)\big)\bigg)^{\frac{2n+1}{2n}}\right]^{\frac{2n}{2n+1}}.$
To prove (4.1), it is enough to show that
$J\lesssim\|f_{1}\|_{L^{\frac{2n+1}{2}}(\mathbb{F}_{q}^{2n},dx)}\|f_{n+1}\|_{L^{\frac{2n+1}{2}}(\mathbb{F}_{q}^{2n},dx)}\|f_{n}\|_{L^{2n+1}(\mathbb{F}_{q}^{2n},dx)}\prod_{j=2}^{n-1}\left(\|f_{j}\|_{L^{2n+1}(\mathbb{F}_{q}^{2n},dx)}\|f_{n+j}\|_{L^{2n+1}(\mathbb{F}_{q}^{2n},dx)}\right).$
Step $2^{\circ}$ (Extract the $f_{n}$-factor). For fixed $y$ and $x_{n}$, perform the change of variables $s:=\tau+x_{n}y$. For $x=(x_{1},\ldots,x_{2n})\in\mathbb{F}_{q}^{2n}$, we use the notation $\widehat{x}_{i,j}$ to denote the element in $\mathbb{F}_{q}^{2n-2}$ obtained by removing $x_{i}$ and $x_{j}$ from $x$. By Minkowski inequality, we obtain
$J=$ $\left[\frac{1}{q^{2n-1}}\sum_{x^{\prime}\in\mathbb{F}_{q}^{2n-1}}\frac{1}{q}\sum_{s\in\mathbb{F}_{q}}\bigg(\frac{1}{q}\sum_{y\in\mathbb{F}_{q}}f_{n}(\widehat{x}_{n},s)\prod_{\begin{subarray}{c}j=1\\ j\neq n,2n\end{subarray}}^{2n}f_{j}\big(\pi_{j}(x^{\prime},y,s-\frac{1}{2}x_{n}y)\big)\bigg)^{\frac{2n+1}{2n}}\right]^{\frac{2n}{2n+1}}$
$\leq$ $\frac{1}{q}\sum_{y\in\mathbb{F}_{q}}\bigg[\frac{1}{q^{2n-1}}\sum_{x^{\prime}\in\mathbb{F}^{2n-1}_{q}}\frac{1}{q}\sum_{s\in\mathbb{F}_{q}}f_{n}(\widehat{x}_{n},s)^{\frac{2n+1}{2n}}\prod_{\begin{subarray}{c}j\neq n,2n\end{subarray}}f_{j}\big(\pi_{j}(x^{\prime},y,s-\frac{1}{2}x_{n}y)\big)^{\frac{2n+1}{2n}}\bigg]^{\frac{2n}{2n+1}}$
$=$ $\frac{1}{q}\sum_{y\in\mathbb{F}_{q}}\bigg[\frac{1}{q^{2n-1}}\sum_{(\widehat{x}_{n,2n},s)\in\mathbb{F}^{2n-1}_{q}}f_{n}(\widehat{x}_{n},s)^{\frac{2n+1}{2n}}\bigg(\frac{1}{q}\sum_{x_{n}\in\mathbb{F}_{q}}\prod_{\begin{subarray}{c}j\neq n,2n\end{subarray}}f_{j}\big(\pi_{j}(x^{\prime},y,s-\frac{1}{2}x_{n}y)\big)^{\frac{2n+1}{2n}}\bigg)\bigg]^{\frac{2n}{2n+1}}$
where in the last equality we have used the fact that $f_{n}$-term is independent of $x_{n}$. By applying Hölder inequality with the conjugate pair of exponents $(2n,\frac{2n}{2n-1})$, we obtain
$J\leq\frac{1}{q}\sum_{y\in\mathbb{F}_{q}}J_{1}(y)J_{2}(y)$
where
$J_{1}(y)=$ $\bigg(\frac{1}{q^{2n-1}}\sum_{(\widehat{x}_{n,2n},s)\in\mathbb{F}^{2n-1}_{q}}f_{n}(\widehat{x}_{n},s)^{2n+1}\bigg)^{\frac{1}{2n+1}},$
$J_{2}(y)=$ $\bigg(\frac{1}{q^{2n-1}}\sum_{(\widehat{x}_{n,2n},s)\in\mathbb{F}_{q}^{2n-1}}\bigg(\frac{1}{q}\sum_{x_{n}\in\mathbb{F}_{q}}\prod_{\begin{subarray}{c}j\neq n,2n\end{subarray}}f_{j}\big(\pi_{j}(x^{\prime},y,s-\frac{1}{2}x_{n}y)^{\frac{2n+1}{2n}}\bigg)^{\frac{2n}{2n-1}}\bigg)^{\frac{2n-1}{2n+1}}$
Applying Hölder inequality in the outer yy–sum with the conjugate pair of exponents $\left(2n+1,\frac{2n+1}{2n}\right)$ yields
$J\lesssim$ $\Big(\frac{1}{q}\sum_{y}J_{1}(y)^{2n+1}\Big)^{\frac{1}{2n+1}}\cdot\Bigg(\frac{1}{q}\sum_{y\in\mathbb{F}_{q}}J_{2}(y)^{\frac{2n+1}{2n}}\Bigg)^{\frac{2n}{2n+1}}$
$=$ $\bigg(\frac{1}{q^{2n}}\sum_{(\widehat{x}_{n},s)\in\mathbb{F}^{2n}_{q}}f_{n}(\widehat{x}_{n},s)^{2n+1}\bigg)^{\frac{1}{2n+1}}\cdot J_{\Pi}$
$=$ $\|f_{n}\|_{L^{2n+1}(\mathbb{F}_{q}^{2n},dx)}\cdot J_{\Pi}.$
Therefore, it remains to bound
$J_{\Pi}:=\Bigg(\frac{1}{q}\sum_{y\in\mathbb{F}_{q}}J_{2}(y)^{\frac{2n+1}{2n}}\Bigg)^{\frac{2n}{2n+1}}$
by the product of the remaining norms appearing on the right-hand side of (4.1).
Step $3^{\circ}$ (Apply the induction hypothesis).
To pass from dimension $n$ to n1n-1, we isolate the last pair of coordinates $(x_{n},x_{2n})$ and treat it as a parameter. For each fixed choice of $(x_{n},x_{2n})\in\mathbb{F}_{q}^{2}$, the remaining variables (together with the $t$-variable after the changes made in Steps $1^{\circ}$$2^{\circ}$) carry the same Heisenberg structure as $\mathbb{H}^{\,n-1}(\mathbb{F}_{q})$. With this preparation, the bracketed expression in (4) can be viewed as an (n1)(n-1)–dimensional Loomis–Whitney type form in the remaining variables, applied to auxiliary factors obtained from the original $f_{j}$’s by freezing $(x_{n},x_{2n})$. This allows us to invoke the induction hypothesis and obtain (4). The specific power $\frac{2n+1}{2n-1}$ is chosen so that the norms produced by the (n1)(n-1)–dimensional estimate translate, after averaging over $(x_{n},x_{2n})$, into the desired $L^{2n+1}$ and $L^{\frac{2n+1}{2}}$ norms of the original functions.
By applying Minkowski’s inequality, we obtain
$J_{\Pi}=$ $\Bigg(\frac{1}{q}\sum_{y\in\mathbb{F}_{q}}J_{2}(y)^{\frac{2n+1}{2n}}\Bigg)^{\frac{2n}{2n+1}}$
$\lesssim$ $\left(\frac{1}{q^{2}}\sum_{x_{n},y\in\mathbb{F}_{q}}\left[\frac{1}{q^{2n-1}}\sum_{(\widehat{x}_{n,2n},t)\in\mathbb{F}_{q}^{2n-1}}\prod_{\begin{subarray}{c}j\neq n,2n\end{subarray}}f_{j}(\pi_{j}(x^{\prime},y,t))^{\frac{2n+1}{2n-1}}\right]^{\frac{2n-1}{2n}}\right)^{\frac{2n}{2n+1}}.$ (19)
Observing that the expression inside the square brackets of (4) is a product of 2n22n-2 functions, we aim to transform it into a form suitable for applying the induction hypothesis to establish an upper bound for $J_{\Pi}$. We recall that
$f_{j}(\pi_{j}(x^{\prime},y,\tau))=\begin{cases}f_{j}(\check{x}_{j},\tau+\frac{1}{2}x_{j}x_{j+n}),~~~~~~~~~~~~~\text{ if }1\leq j\leq n-1,\\ f_{j}(\check{x}_{n+j},\tau-\frac{1}{2}x_{j-n}x_{j}),~~~\text{ if }n+1\leq j\leq 2n-1.\end{cases}$ (20)
We temporarily represent points in $\mathbb{H}^{n-1}(\mathbb{F}_{q})$ using coordinates $(u,t)=(u_{1},\ldots,u_{2n-2},t)$ where $u\in\mathbb{F}_{q}^{2n-2}$. Fixing $x_{n},y\in\mathbb{F}_{q}$ we define the functions $g_{j}$, $j\in\{1,\ldots,2n-2\}$, on $\mathbb{F}_{q}^{2n-2}$ by
$g_{j}(\widehat{u}_{j},t)=\begin{cases}f_{j}(\tilde{u}_{j})^{\tfrac{2n+1}{2n-1}},~~~~~~~~~~1\leq j\leq n-1,\\ f_{j+1}(\tilde{u}_{j})^{\tfrac{2n+1}{2n-1}},~~~~~n\leq j\leq 2n-2,\end{cases}$
where $\tilde{u}_{j}$ is obtained from
$(u_{1},u_{2},\ldots,u_{n-1},x_{n},u_{n},\ldots,u_{2n-2},y,t)$
by removing $u_{j}$. More precisely
$g_{j}(\widehat{u}_{j},t)=\begin{cases}f_{1}(u_{2},\ldots,u_{n-1},x_{n},u_{n},\ldots,u_{2n-2},y,t)^{\frac{2n+1}{2n-1}},~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~j=1,\\ f_{j}(u_{1},\ldots,u_{j-1},u_{j+1},\ldots,u_{n-1},x_{n},u_{n},\ldots,u_{2n-2},y,t)^{\frac{2n+1}{2n-1}},~~~2\leq j\leq n-2,\\ f_{n-1}(u_{1},\ldots,u_{n-2},x_{n},u_{n},\ldots,u_{2n-2},y,t)^{\frac{2n+1}{2n-1}},~~~~~~~~~~~~~~~~~~~~~~~~~j=n-1,\end{cases}$ (21)
and
$g_{j}(\widehat{u}_{j},t)=\begin{cases}f_{n+1}(u_{1},\ldots,u_{n-1},x_{n},u_{n+1},\ldots,u_{2n-2},y,t)^{\frac{2n+1}{2n-1}},~~~~~~~~~~~~~~~~~~~~~~~~~j=n,\\ f_{j}(u_{1},\ldots,u_{n-1},x_{n},u_{n},\ldots,u_{j-1},u_{j+1},\ldots,u_{2n-2},y,t)^{\frac{2n+1}{2n-1}},~n+1\leq j\leq 2n-3,\\ f_{2n-1}(u_{1},\ldots,u_{n-2},x_{n},u_{n},\ldots,u_{2n-3},y,t)^{\frac{2n+1}{2n-1}},~~~~~~~~~~~~~~~~~~~~~~~~~j=2n-2.\end{cases}$ (22)
With this notation in place, and recalling (20), we can restate (4) equivalently as follows
$J_{\Pi}\lesssim$ $\left(\frac{1}{q^{2}}\sum_{x_{n},y\in\mathbb{F}_{q}}\left[\frac{1}{q^{2n-1}}\sum_{(u,t)\in\mathbb{F}_{q}^{2n-1}}\prod_{\begin{subarray}{c}1\leq j\leq 2n-2\end{subarray}}g_{j}(\pi_{j}(u,t))\right]^{\frac{2n-1}{2n}}\right)^{\frac{2n}{2n+1}}$
where we recall that $\pi_{j}$ is the projection from $\mathbb{H}^{n-1}(\mathbb{F}_{q})$ to the vertical plane $W_{j}$ (identified with $\mathbb{F}_{q}^{2n-2}$), defined in (5). By the induction hypothesis we have
$\frac{1}{q^{2n-1}}$ $\sum_{(u,t)\in\mathbb{F}_{q}^{2n-1}}\prod_{\begin{subarray}{c}1\leq j\leq 2n-2\end{subarray}}g_{j}(\pi_{j}(u,t))$
$\;\lesssim\;$ $\|g_{1}\|_{L^{\frac{2n-1}{2}}(\mathbb{F}_{q}^{2n-2},dx)}\,\|g_{n}\|_{L^{\frac{2n-1}{2}}(\mathbb{F}_{q}^{2n-2},dx)}\,\prod_{\begin{subarray}{c}1\leq j\leq 2n-2,\\ j\neq 1,n\end{subarray}}\|g_{j}\|_{L^{2n-1}(\mathbb{F}_{q}^{2n-2},dx)}.$ (23)
Next we apply the multilinear Hölder’s inequality with exponents $\ell_{1}=\ell_{n}=n$ and $\ell_{j}=2n$, $\forall\,j\neq 1,n$. Note that $\sum\limits_{j=1}^{2n-2}\frac{1}{\ell_{j}}=1$. Hence, from (4), we have
$J_{\Pi}\lesssim$ $\bigg[\bigg(\frac{1}{q^{2}}\sum_{x_{n},y\in\mathbb{F}_{q}}\|g_{1}\|_{L^{\frac{2n-1}{2}}(\mathbb{F}_{q}^{2n-2},dx)}^{\frac{2n-1}{2}}\bigg)^{\frac{1}{n}}\cdot\bigg(\frac{1}{q^{2}}\sum_{x_{n},y\in\mathbb{F}_{q}}\|g_{n}\|_{L^{\frac{2n-1}{2}}(\mathbb{F}_{q}^{2n-2},dx)}^{\frac{2n-1}{2}}\bigg)^{\frac{1}{n}}$
$\cdot\prod_{\begin{subarray}{c}1\leq j\leq 2n-2,\\ j\neq 1,n\end{subarray}}\bigg(\frac{1}{q^{2}}\sum_{x_{n},y\in\mathbb{F}_{q}}\|g_{j}\|_{L^{2n-1}(\mathbb{F}_{q}^{2n-2},dx)}^{2n-1}\bigg)^{\frac{1}{2n}}\bigg]^{\frac{2n}{2n+1}}$
$=$ $\|f_{1}\|_{L^{\frac{2n+1}{2}}(\mathbb{F}_{q}^{2n},dx)}\|f_{n+1}\|_{L^{\frac{2n+1}{2}}(\mathbb{F}_{q}^{2n},dx)}\prod_{j=2}^{n-1}\left(\|f_{j}\|_{L^{2n+1}(\mathbb{F}_{q}^{2n},dx)}\|f_{n+j}\|_{L^{2n+1}(\mathbb{F}_{q}^{2n},dx)}\right).$
Here the last equality follows from the definition of $g_{j}$ for $j\in\{1,\ldots,2n-2\}$, given in (21) and (22).
For a positive $n$, we define a multilinear map
$T:\mathscr{F}(\mathbb{F}_{q}^{2n})\times\cdots\times\mathscr{F}(\mathbb{F}_{q}^{2n})\rightarrow\mathscr{F}(\mathbb{F}_{q}^{2n})$
as follows: For an element $(z,\tau)=(z_{1},\ldots,z_{2n-1},\tau)\in\mathbb{F}_{q}^{2n}$
$T(g_{1},\ldots,g_{2n-1})(z,\tau):=\frac{1}{q}\sum_{s\in\mathbb{F}_{q}}\left[g_{n}\big(z,\tau+z_{n}s\big)\cdot\prod_{\begin{subarray}{c}j=1,j\neq n\end{subarray}}^{2n-1}g_{j}\left(\pi_{j}\big(z,s,\tau+\frac{1}{2}z_{n}s\big)\right)\right].$
We use this operator $T$ to find an upper bound of $I$, given in (4). More precisely, taking a change of variable $\tau=t+\frac{1}{2}x_{n}x_{2n}$, and applying Hölder’s inequality with the conjugate pair of exponents
$(\ell,\ell^{\prime})=\bigg(\frac{n(2n+1)}{n+1},\frac{n(2n+1)}{2n^{2}-1}\bigg),$ (24)
we obtain
$I=$ $\frac{1}{q^{2n+1}}\sum_{(x,t)\in\mathbb{H}^{n}(\mathbb{F}_{q})}\prod_{j=1}^{2n}f_{j}(\pi_{j}(x,t))$
$=$ $\frac{1}{q^{2n}}\sum_{(x^{\prime},t)\in\mathbb{F}_{q}^{2n}}\sum_{x_{2n}\in\mathbb{F}_{q}}f_{2n}(x^{\prime},t-\frac{1}{2}x_{n}x_{2n})\,\Bigg(\frac{1}{q}f_{n}(\widehat{x}_{n},t+\frac{1}{2}x_{n}x_{2n})\prod_{\begin{subarray}{c}j\neq n,2n\end{subarray}}f_{j}\big(\pi_{j}(x,t)\big)\Bigg)$
$=$ $\frac{1}{q^{2n}}\sum_{(x^{\prime},\tau)\in\mathbb{F}_{q}^{2n}}f_{2n}(x^{\prime},\tau)\,\Bigg(\frac{1}{q}\sum_{x_{2n}\in\mathbb{F}_{q}}f_{n}(\widehat{x}_{n},\tau-\frac{1}{2}x_{n}x_{2n})\prod_{\begin{subarray}{c}j\neq n,2n\end{subarray}}f_{j}\big(\pi_{j}(x^{\prime},x_{2n},\tau+\frac{1}{2}x_{n}x_{2n})\big)\Bigg)$
$\leq$ $\|T(f_{1},\ldots,f_{2n-1})\|_{L^{\ell^{\prime}}(\mathbb{F}^{2n}_{q},dx)}\cdot\|f_{2n}\|_{L^{\ell}(\mathbb{F}^{2n}_{q},dx)}.$ (25)
We are now ready to prove Theorem 1.4.
Proof of Theorem 1.4.
In view of (4), the proof of Theorem 1.4 is reduced to proving
$\|T(f_{1},\ldots,f_{2n-1})\|_{L^{\ell^{\prime}}(\mathbb{F}^{2n}_{q},dx)}\lesssim\prod_{j=1}^{2n-1}\|f_{j}\|_{L^{\ell_{j}}(\mathbb{F}^{2n}_{q},dx)}$ (26)
for all nonnegative $f_{1},\ldots,f_{2n-1}\in\mathscr{\mathbb{F}_{q}^{2n}}$, where $\ell^{\prime}$ is defined in (24) and $\ell_{1}=\ldots=\ell_{2n-1}=\frac{n(2n+1)}{n+1}$. To prove the inequality (26), first, from (4.1) we obtain
$\|T(f_{1},\ldots,f_{2n-1})\|_{L^{\ell_{k}}(\mathbb{F}^{2n}_{q},dx)}\lesssim\prod_{j=1}^{2n-1}\|f_{j}\|_{L^{\ell_{j,k}}(\mathbb{F}^{2n}_{q},dx)},~~~k=1,\ldots,n$ (27)
where $\ell_{k}$, $\ell_{j,k}$ are given by
$\ell_{k}=\begin{cases}\frac{2n+1}{2n},~~~~~~~~~~k=1,\ldots,n-1,\\ \frac{2n+1}{2n-1},~~~~~~~~~~k=n\end{cases}$ (28)
and
[…]
$j=\{1,\ldots,2n-1\}$, $k=1,\ldots,n$.
Note that the exponents in (28) and (29) satisfy
$\frac{1}{\ell^{\prime}}=\frac{2n^{2}-1}{n(2n+1)}=\frac{1}{n}\bigg[\frac{2n}{2n+1}\cdot(n-1)+\frac{2n-1}{2n}\bigg]=\sum\limits_{k=1}^{n}\frac{1}{n}\frac{1}{\ell_{k}}$
and
$\frac{1}{\ell_{j}}=\sum\limits_{k=1}^{n}\frac{1}{n}\frac{1}{\ell_{j,k}},~~~~~~~~j=1,\ldots,2n-1.$
To prove (26), we apply the discrete version of the multilinear interpolation theorem, a generalization to [14, Proposition 6.1]. Since $(\mathbb{F}_{q}^{2n},dx)$ is a finite measure space equipped with the normalized counting measure, the same interpolation argument carries over verbatim. We then perform (n1)(n-1) iterations with $\theta=\tfrac{1}{k}$, starting from $k=2$ and ending at $k=n$, using the identity
$\frac{1}{k}[a_{1}+\cdots+a_{k}]=\left(1-\frac{1}{k}\right)\left(\frac{1}{k-1}[a_{1}+\cdots+a_{k-1}]+\frac{1}{k}a_{k}\right).$
More precisely, we first apply multilinear interpolation theorem with $\theta=\frac{1}{2}$ to the two operator bounds given by (27) for $k=1$ and $k=2$. Next, we apply the multilinear interpolation theorem with $\theta=\frac{1}{3}$ to interpolate between the newly obtained bound and the operator bound stated in (27) for $k=3$. We continue this process until, in the final step, we apply the theorem with $\theta=\frac{1}{n}$ to interpolate between the previously obtained bound and the bound for $k=n$. From this, we obtain (26) for all nonnegative $f_{1},\ldots,f_{2n}\in\mathscr{F}(\mathbb{F}^{2n}_{q})$.
5 Proof of Corollary 1.5
Proof.
For subsets $K_{1},\ldots,K_{2n}\subset\mathbb{F}_{q}^{2n}$, denote
$K:=\bigcap\limits_{j=1}^{2n}\pi_{j}^{-1}(K_{j}),$
where each $K_{j}$ is viewed as a subset of $\mathbb{H}^{n}(\mathbb{F}_{q})$ through the identification $\mathbb{F}_{q}^{2n}=W_{j}\subset\mathbb{H}^{n}(\mathbb{F}_{q})$ for $j\in\{1,\ldots 2n\}$. Then for $(x,t)\in\mathbb{H}^{n}(\mathbb{F}_{q}),$ we can write
$\mathbf{1}_{K}(x,t)=\prod_{j=1}^{2n}\mathbf{1}_{K_{j}}(\pi_{j}(x,t)).$
Then, to prove Corollary 1.5, it is sufficient to prove the inequality
$\frac{1}{q^{2n+1}}\sum_{(x,t)\in\mathbb{H}^{n}(\mathbb{F}_{q})}\prod_{j=1}^{2n}\mathbf{1}_{K_{j}}(\pi_{j}(x,t))\;\;\lesssim\;\;\prod_{j=1}^{2n}\|\mathbf{1}_{K_{j}}\|_{L^{\frac{n(2n+1)}{n+1}}(\mathbb{F}_{q}^{2n},dx)}.$ (30)
Note that
$\|\mathbf{1}_{K_{j}}\|_{L^{p}}=\bigg(\frac{|K_{j}|}{q^{2n}}\bigg)^{1/p},\qquad j=1,\dots,2n.$
Finally, applying Theorem 1.4 with $f_{j}=\mathbf{1}_{K_{j}}$ for $j\in\{1,\ldots,2n\}$, multiply both sides by $q^{2n+1}$ and simplify the power of $q$ to obtain (30), as desired. ∎
6 Proof of Theorem 1.6
Let $\mathscr{P}$ be a set of points in $\mathbb{F}_{q}^{2}$ and $\mathscr{L}$ be a set of lines in $\mathbb{F}_{q}^{2}$. The number of incidences between $\mathscr{P}$ and $\mathscr{L}$, denoted by $I(\mathscr{P},\mathscr{L})$, is defined by
$I(\mathscr{P},\mathscr{L})=|\{(p,\ell)\in\mathscr{P}\times\mathscr{L}\colon p\in\ell\}|.$
To prove Theorem 1.6, we invoke the following incidence bound due to Vinh [20].
Theorem 6.1.
Let $\mathscr{P}$ be a set of points in $\mathbb{F}_{q}^{2}$ and $\mathscr{L}$ a set of lines in $\mathbb{F}_{q}^{2}$. The number of incidences between $\mathscr{P}$ and $\mathscr{L}$ satisfies
$I(\mathscr{P},\mathscr{L})\leq\frac{|\mathscr{P}|\,|\mathscr{L}|}{q}+2q^{\frac{1}{2}}\sqrt{|\mathscr{P}|\,|\mathscr{L}|}.$
Proof of Theorem 1.6.
As in Section 1, we define a bijective map $\phi\colon\mathbb{F}_{q}^{3}\to\mathbb{F}_{q}^{3}$ by $\phi(x,y,t)=(x,y,t+\frac{xy}{2})$, then
$\frac{1}{q^{3}}\sum_{(x_{1},x_{2},t)\in\mathbb{H}_{q}^{1}}f_{1}(\pi_{1}(x_{1},x_{2},t))f_{2}(\pi_{2}(x_{1},x_{2},t))=\frac{1}{q^{3}}\sum_{x_{1},x_{2},t\in\mathbb{F}_{q}}f_{1}(\pi_{1}(\phi(x_{1},x_{2},t)))f_{2}(\pi_{2}(\phi(x_{1},x_{2},t)))$
$=\frac{1}{q^{3}}\sum_{x_{1},x_{2},t\in\mathbb{F}_{q}}f_{1}(x_{2},t+x_{1}x_{2})f_{2}(x_{1},t)$
$=\frac{1}{q^{3}}\sum_{(a_{1},a_{2}),(b_{1},b_{2})\in\mathbb{F}^{2}_{q},a_{1}b_{1}+b_{2}=a_{2}}f_{1}(a_{1},a_{2})f_{2}(b_{1},b_{2}).$
Let $f_{1}$ and $f_{2}$ be the indicator functions of $\pi_{1}(K)$ and $\pi_{2}(K)$, respectively. Each element $(x_{1},x_{2},t)\in K$ determines a point $(a_{1},a_{2})=\pi_{1}(x_{1},x_{2},t)\in P$ and a line $\ell=\pi_{2}(x_{1},x_{2},t)\in L$, which satisfy the relation $a_{1}b_{1}+b_{2}=a_{2}$. Hence each element of $K$ gives rise to an incidence between $P$ and $L$.
Moreover, when $n=1$, the map
$(x_{1},x_{2},t)\longmapsto\bigl(\pi_{1}(x_{1},x_{2},t),\,\pi_{2}(x_{1},x_{2},t)\bigr)$
is injective, so distinct elements of $K$ yield distinct incidences. Therefore,
$|K|\leq I(P,L).$
Applying Theorem 6.1, we obtain
$\frac{|K|}{q^{3}}\leq\frac{|\pi_{1}(K)|\,|\pi_{2}(K)|}{q^{4}}+2q^{-5/2}\sqrt{|\pi_{1}(K)|\,|\pi_{2}(K)|}.$
This yields the desired conclusion. ∎
7 Connections to other topics
In this section, we address how our results are connected to some questions in the literature.
7.1 Bhowmik, Iosevich, Koh, and Pham’s boundedness problem
Let $\mathcal{K}\colon\mathbb{F}_{q}^{d}\times\mathbb{F}_{q}^{d}\to\mathbb{R}$ be any function (which, below, will play the role of the kernel of an integral operator for the counting measure), and $\mathcal{G}=(V,E)$ a connected ordered graph on $m$ vertices. In a recent paper [4], Bhowmik, Iosevich, Koh, and Pham introduced the boundedness problem for a multilinear operator associated with $\mathcal{K}$ and $\mathcal{G}$. To be precise, define
$\Lambda=\Lambda_{\mathcal{G}}^{\mathcal{K}}:\mathscr{F}(\mathbb{F}_{q}^{d})\times\cdots\times\mathscr{F}(\mathbb{F}_{q}^{d})\rightarrow\mathbb{R}$
by
$(f_{1},\ldots,f_{m})\mapsto\frac{1}{\mathcal{N}(\mathcal{G})}\sum_{x^{1},\dots,x^{m}\in{\mathbb{F}}_{q}^{d}}\ \prod_{(i,j)\in{E}(\mathcal{G})}\mathcal{K}(x^{i},x^{j})\prod_{l=1}^{m}f_{l}(x^{l}),$
where $\mathcal{N}(\mathcal{G})$ is the number of distinct embeddings of $\mathcal{G}$ in $\mathbb{F}_{q}^{d}$. The boundedness problem for operator $\Lambda_{\mathcal{G}}^{\mathcal{K}}$ asks if one can determine numbers $1\leq u_{1},\ldots,u_{m}\leq\infty$ for which there is a (smallest) constant $\Lambda_{\mathcal{G}}^{\mathcal{K}}(u_{1},\dots,u_{m})>0$ such that the inequality
$\frac{1}{\mathcal{N}(\mathcal{G})}\sum_{x^{1},\dots,x^{m}\in{\mathbb{F}}_{q}^{d}}\ \prod_{(i,j)\in{E}(\mathcal{G})}\mathcal{K}(x^{i},x^{j})\prod_{l=1}^{m}f_{l}(x^{l})\leq\Lambda_{\mathcal{G}}^{\mathcal{K}}(u_{1},\dots,u_{m})\prod_{i=1}^{m}{\|f_{i}\|}_{L^{u_{i}}(\mathbb{F}_{q}^{d},dx)}$ (31)
holds for all nonnegative $f_{1},\ldots,f_{m}\in\mathscr{F}(\mathbb{F}_{q}^{d})$. Using the discrete Fourier analysis, they settled the case where $\mathcal{G}$ is a connected graph with at most $4$ vertices and $\mathcal{K}(x,y)$ is defined by $\mathcal{K}(x,y)=1$ if $\|x-y\|=(x_{1}-y_{1})^{2}+(x_{2}-y_{2})^{2}=1$, and $\mathcal{K}(x,y)=0$ otherwise.
The operator $\Lambda_{\mathcal{G}}^{\mathcal{K}}$ is a generalization of several geometric counting functions:
  1. 1.
    If $\mathcal{G}=K_{2}$, and $f_{1}=f_{2}$ are the indicator functions of a set $A$ in $\mathbb{F}_{q}^{d}$, then the operator $\Lambda_{\mathcal{G}}^{\mathcal{K}}$ counts the number of pairs of points in $A\times A$ of distance $1$.
  2. 2.
    If $\mathcal{G}=C_{4}$, and $f_{1}=f_{2}=f_{3}=f_{4}$ are the indicator functions of a set $A$ in $\mathbb{F}_{q}^{d}$, then the operator $\Lambda_{\mathcal{G}}^{\mathcal{K}}$ counts the number of rhombi of side-length $1$ in $A$.
  3. 3.
    If $\mathcal{G}$ is a tree and $f_{1}=\cdots=f_{m}$ are the indicator functions of a set $A$ in $\mathbb{F}_{q}^{d}$, then the operator $\Lambda_{\mathcal{G}}^{\mathcal{K}}$ counts the number of copies of the tree $\mathcal{G}$ in $A$.
The inequality (7) for $n=1$ is essentially the same as the inequality (31) for $m=2$, with suitable $\mathcal{G}$ and $\mathcal{K}$ chosen. To see this, define $\phi\colon\mathbb{F}_{q}^{3}\to\mathbb{F}_{q}^{3}$ by $\phi(x,y,t)=(x,y,t+\frac{xy}{2})$. Then, since $\phi$ is a bijective map, the left-hand side of (7) can be written
$\frac{1}{q^{3}}\sum_{(x_{1},x_{2},t)\in\mathbb{H}_{q}^{1}}f_{1}(\pi_{1}(x_{1},x_{2},t))f_{2}(\pi_{2}(x_{1},x_{2},t))=\frac{1}{q^{3}}\sum_{x_{1},x_{2},t\in\mathbb{F}_{q}}f_{1}(\pi_{1}(\phi(x_{1},x_{2},t)))f_{2}(\pi_{2}(\phi(x_{1},x_{2},t)))$
$=\frac{1}{q^{3}}\sum_{x_{1},x_{2},t\in\mathbb{F}_{q}}f_{1}(x_{2},t+x_{1}x_{2})f_{2}(x_{1},t).$ (32)
Now let $\mathcal{G}$ be the complete graph of two vertices and define $\mathcal{K}$ by
$\mathcal{K}((a_{1},a_{2}),(b_{1},b_{2}))=\begin{cases}1~&\mbox{if}~a_{1}b_{1}+b_{2}-a_{2}=0,\\ 0~&\mbox{otherwise}.\end{cases}$
Then the left-hand side $\Lambda_{\mathcal{G}}^{\mathcal{K}}(f_{1},f_{2})$ of the inequality (31) is equal to
$\frac{1}{q^{3}}\sum_{(a_{1},a_{2}),(b_{1},b_{2})\in\mathbb{F}^{2}_{q},a_{1}b_{1}+b_{2}=a_{2}}f_{1}(a_{1},a_{2})f_{2}(b_{1},b_{2}),$
which coincides with the value (7.1). Thus, in this case, both inequalities (7) and (31) coincide with each other. This connection suggests a further study of the operator $\Lambda_{\mathcal{G}}^{\mathcal{K}}$ for various $\mathcal{K}$ and $\mathcal{G}$.
7.2 The orthogonal projection problem
For a subset $K\subset\mathbb{F}_{q}^{2n+1}$ and an integer $r\geq 1$, let
$\mathscr{E}_{r}=\mathscr{E}_{r}(K)\;:=$ $\;\big\{\,W\leq\mathbb{F}_{q}^{2n+1}:\ \dim W=2n,\;\;K\text{ can be covered by at most }r$
 $~~~~\text{ (additive) translates of }W^{\perp}\,\big\},$
where a translate of $W^{\perp}$ means a subset of $\mathbb{F}_{q}^{2n+1}$ of the form $(x,t)+W^{\perp}$ for some $(u,\tau)\in\mathbb{F}_{q}^{2n+1}$ Chen [6] proved that
$|\mathscr{E}_{r}|\;\leq\;\begin{cases}q^{2n-1}\,r,&\text{if }r\leq\dfrac{|K|}{2},\\[5.69054pt] \dfrac{r\,q^{4n}}{(q^{2n}-r)\,|K|},&\text{if }0<r<q^{2n}.\end{cases}$ (33)
Informally, this states that for most subspaces $W\subset\mathbb{F}_{q}^{2n+1}$, the covering number of $K$ by translates of $W^{\perp}$ is not very small.
Chen’s result can be interpreted from the view of the (orthogonal) projection problem. To see this, for each $j=1,\ldots,2n$, let $P_{j}:\mathbb{F}_{q}^{2n+1}\mapsto W_{j}$ be the orthogonal projection defined by $P_{j}(x,t):=(\check{x}_{j},t),$ where $W_{j}$ is the coordinate hyperplane given in (3). Note that $W_{j}^{\perp}=\mathrm{Ker}\,P_{j}$ and the translates of $W_{j}^{\perp}$ are exactly cosets of the subspace $W_{j}^{\perp}$ in $\mathbb{F}_{q}^{2n+1},$ which are in turn the fibres of $P_{j}$. Note that each translate of $W_{j}^{\perp}$ can be written in the form $(u,\tau)+W_{j}^{\perp}$ for uniquely determined $(u,\tau)\in W_{j}.$ Thus, we observe that for a subset $K$ of $\mathbb{F}_{q}^{2n+1}$, $K\cap((u,\tau)+W_{j}^{\perp})\neq\emptyset$ for $(u,\tau)\in W_{j}$ if and only if $(u,\tau)\in P_{j}(K)$. Thus, one can write
$K\;\subset\;\bigcup_{(u,\tau)\in P_{j}(K)}\left((u,\tau)+W_{j}^{\perp}\right).$ (34)
This implies that the minimal number of translates of $W_{j}^{\perp}$ that cover a set $K\subset\mathbb{F}_{q}^{2n+1}$ coincides with $|P_{j}(K)|.$
Now let $\mathscr{H}=\{W_{1},\dots,W_{2n}\}$ be the set of coordinate hyperplanes in $\mathbb{F}_{q}^{2n+1}$. If $|\mathscr{H}|=2n>|\mathscr{E}_{r}|$, then, by the definition of $\mathscr{E}_{r}$, there exists at least one coordinate hyperplane $W_{j}\in\mathscr{H}$ that does not belong to $\mathscr{E}_{r}$, and for this $j$ the covering number of $K$ by translates of $W_{j}^{\perp}$ is strictly larger than $r$, i.e., $P_{j}(K)$ contains more than $r$ points.
Combining this observation with Chen’s estimate (33) and the fact that $|\mathscr{H}|=2n$, one finds a constant $c_{n}>0$ depending only on $n$ such that for every nonempty set $K\subset\mathbb{F}_{q}^{2n+1}$ there exists an index $j_{0}\in\{1,\dots,2n\}$ with
$|P_{j_{0}}(K)|\;\geq\;c_{n}\,\frac{|K|}{q^{2n}}.$ (35)
Thus, among the given $2n$ projections $P_{j}:\mathbb{F}_{q}^{2n+1}\mapsto W_{j}$ for $j=1,\ldots,2n,$ there exists at least one projection $P_{j_{0}}$ under which the image of $K$ must have the size comparable to $|K|/q^{2n}$.
This discussion can be phrased in a more general framework. More precisely, let $(H,\cdot)$ be a (finite) set equipped with an algebraic structure with the identity and let $\{\Phi_{i}:H\mapsto V_{i}\,|\,i\in I\}$ be a family of algebraic projections, where $V_{i}$ are subobjects of $H$ with the same cardinality. Let $Z_{i}:=Ker\,\Phi_{i}$. Then the situation is analogous to the case of the orthogonal projections $P_{j}$: Each coset of $Z_{j}$ can be expressed in the form $v\cdot Z_{j}$ for uniquely determined $v\in V_{j},$ and a coset $v\cdot Z_{j}$ for $v\in V_{j}$ is equal to the fiber $\Phi_{i}^{-1}(v).$ Also we have that for a subset $K$ of $H$, $K\cap v\cdot Z_{i}\neq\emptyset$ for $v\in V_{i}$ if and only if $v\in\Phi_{i}(K)$. Thus, one can write
$K\;\subset\;\bigcup_{v\in\Phi_{i}(K)}v\cdot Z.$ (36)
In particular, the minimal number of cosets of $Z_{i}$ in $H$ that cover $K$ is equal to $|\Phi_{i}(K)|.$ In the light of (35), one seek an index $i_{0}\in I$ for which the image $\Phi_{i_{0}}(K)\subset V_{i_{0}}$ is as large as possible. In practice, depending on which features of the original set one wishes to preserve, the choice of an appropriate family of projection maps is crucial. Our case concerns $H=\mathbb{H}^{n}(\mathbb{F}_{q})$ and $\Phi_{i}=\pi_{i}$ for $i\in I=\{1,\ldots,2n\}.$
In the following, we compare Chen’s theorem and Corollary 1.5 from this point of view. For each $j=1,\ldots,2n$, define $T_{j}:\mathbb{H}^{n}(\mathbb{F}_{q})\to\mathbb{F}_{q}^{2n+1}$ by
$T_{j}(x,t)=\begin{cases}(x,\ t+\tfrac{1}{2}\,x_{j}x_{n+j}),&1\leq j\leq n,\\[2.0pt] (x,\ t-\tfrac{1}{2}\,x_{j-n}\,x_{j}),&n+1\leq j\leq 2n.\end{cases}$ (37)
This map is a bijective map with $P_{j}\circ T_{j}=\pi_{i}$ for all $j=1,\ldots,2n$. We observe that for each $(u,\tau)\in W_{j}$, we have
$T_{j}\big(\pi_{j}^{-1}(u,\tau)\big)=P_{j}^{-1}(u,\tau),$ (38)
i.e., $T_{j}$ maps each Heisenberg fibre $\pi^{-1}(u,\tau)$ bijectively onto the fibre $P_{j}^{-1}(u,\tau)$. Moreover, the restriction of $T_{j}$ to $W_{j}$ is the identity map.
Let $K\subset\mathbb{\mathbb{}}H^{n}(\mathbb{F}_{q})$ and fix $j=1,\ldots,2n.$ Recall that $\mathrm{Ker}\,\pi_{j}=W_{j}^{\perp}$. Letting $H=\mathbb{H}^{n}(\mathbb{F}_{q})$ and $\Phi_{j}=\pi_{j}$, by (36) we have
$K\;\subset\;\bigcup_{(u,\tau)\in\pi_{j}(K)}(u,\tau)\cdot W_{j}^{\perp}.$ (39)
Applying $T_{j}$ on both sides of (39), by the equality (38) and the facts that $\pi_{j}^{-1}(u,\tau)=(u,\tau)\cdot W_{j}^{\perp}$ and $P_{j}^{-1}(u,\tau)=(u,\tau)+W_{j}^{-1}$, we obtain
$T_{j}(K)\;\subset\;\bigcup_{(u,\tau)\in\pi_{j}(K)}\bigl((u,\tau)+W_{j}^{\perp}\bigr).$
Note that each coset $(u,\tau)+W_{j}^{\perp}$ in the above union is necessary to cover $T_{j}(K).$ Thus, the quantity $|\pi_{j}(K)|$ coincides with the covering number of $T_{j}(K)$ by translates of $W_{j}^{\perp}$. Now for each $j=1,\ldots,2n$, define
$\phi_{j}\;:=\;T_{j}^{-1}\circ P_{j}\circ T_{j}:\mathbb{H}^{n}(\mathbb{F}_{q})(\cong\mathbb{F}_{q}^{2n+1})\to\mathbb{H}^{n}(\mathbb{F}_{q})(\cong\mathbb{F}_{q}^{2n+1}).$
These maps enjoy three basic features that we use repeatedly: they are idempotent retractions,
$\phi_{j}\circ\phi_{j}=\phi_{j};$
they have the same fibres as $\pi_{j}$, in the sense that $\phi_{j}(x)=\phi_{j}(y)$ if and only if $\pi_{j}(x)=\pi_{j}(y)$; and, consequently, for every subset $K\subset\mathbb{H}^{n}(\mathbb{F}_{q})$,
$|\phi_{j}(K)|=|\pi_{j}(K)|=\bigl|P_{j}\bigl(T_{j}(K)\bigr)\bigr|.$
Corollary 1.5 yields a constant $C_{n}>0$ such that
$\max_{1\leq j\leq 2n}|\pi_{j}(K)|\;\geq\;C_{n}\,|K|^{\frac{2n+1}{2(n+1)}}q^{-\frac{1}{2(n+1)}},$
and hence there exists an index $j^{\prime}\in\{1,\dots,2n\}$ for which
$|\phi_{j^{\prime}}(K)|\;=\;|P_{j^{\prime}}(T_{j^{\prime}}(K))|\;\geq\;C_{n}\,|K|^{\frac{2n+1}{2(n+1)}}q^{-\frac{1}{2(n+1)}}.$ (40)
We now compare between (35) and (40). Using only the trivial constraint $|K|\leq q^{2n+1}$, one checks that
$\frac{|K|^{\frac{2n+1}{2(n+1)}}q^{-\frac{1}{2(n+1)}}}{|K|/q^{2n}}\;\geq\;q^{2n-1}$
for all nonempty $K\subset\mathbb{H}^{n}(\mathbb{F}_{q})(\cong\mathbb{F}_{q}^{2n+1})$.
Thus, with the same $\mathscr{H}$, the family $\{\pi_{j}\}_{j}$ of projections provides a strictly stronger lower bound for the maximal projection size in one of the vertical directions.
Another related question is a discrete analogue to the problem considered by Balogh, Fässler, Mattila, and Tyson [2] for the real Heisenberg group. To explain this, for $K\subset\mathbb{H}^{n}(\mathbb{F}_{q})$ and $1\leq m\leq n$, let $\mathscr{E}_{r}^{h,m}(K)$ denote the set of homogeneous subgroups $G\subset\mathbb{H}^{n}(\mathbb{F}_{q})$ of size $q^{m}$ such that $K$ can be covered by at most $r$ left cosets of $G^{\perp}$. (See Section 2 for the notion homogeneous and the notation $G^{\perp}$.) Then following Balogh, Fässler, Mattila, and Tyson, it is natural to seek upper bounds for $|\mathscr{E}_{r}^{h,m}(K)|$ in terms of $|K|$, $q$, $n$, and $m$. We plan to pursue this Heisenberg covering problem in a sequel paper.
8 Acknowledgements
Thang Pham and Dung The Tran would like to thank VIASM for the hospitality and for the excellent working conditions. Daewoong Cheong sincerely thanks the Korea Institute for Advanced Study for financial support and hospitality and was partially supported by the research year program of Chungbuk National University in 2025.